### v1

Abstract
Large-scale pre-trained models have achieved remarkable success in language and image tasks, leading an increasing number of studies to explore the application of pre-trained image models, such as CLIP, in the domain of few-shot action recognition (FSAR). However, current methods generally suffer from several problems: 1) Direct fine-tuning often undermines the generalization capability of the pre-trained model; 2) The exploration of task-specific information is insufficient in the visual tasks; 3) The semantic order information is typically overlooked during text modeling; 4) Existing cross-modal alignment techniques ignore the temporal coupling of multimodal information. To address these, we propose Task-Adapter++, a parameter-efficient dual adaptation method for both image and text encoders. Specifically, to make full use of the variations across different few-shot learning tasks, we design a task-specific adaptation for the image encoder so that the most discriminative information can be well noticed during feature extraction. Furthermore, we leverage large language models (LLMs) to generate detailed sequential sub-action descriptions for each action class, and introduce semantic order adapters into the text encoder to effectively model the sequential relationships between these sub-actions. Finally, we develop an innovative fine-grained cross-modal alignment strategy that actively maps visual features to reside in the same temporal stage as semantic descriptions. Extensive experiments fully demonstrate the effectiveness and superiority of the proposed method, which achieves state-of-the-art performance on 5 benchmarks consistently. The code is open-sourced at https://github.com/Jaulin-Bage/Task-Adapter-pp.
keywords:
[…]
Figure 1: Comparison of different paradigms. (a) Existing methods fully fine-tune the feature extractor and combine it with temporal alignment module. (b) We froze the image encoder and apply Task-Specific adaptation to extract the most discriminative information. (c) We use semactic order adapter to model sequential information within sub-actions and further calculate the score by stage with aligned visual features.
1 Introduction
With the advent of large-scale pre-trained models [31, 19], researchers are now able to exploit the knowledge these models have acquired during pre-training and apply it to downstream tasks. Accordingly, there is an increasing focus on fine-tuning large pre-trained models for specific tasks [10, 63, 54, 49]. In video understanding, existing works typically fine-tune pre-trained models on well-labeled video datasets [5, 11], where the training and testing categories are identical. However, in realistic scenarios, collecting sufficient and well-labeled video samples is inherently challenging. As a result, few-shot learning has garnered increasing attention [29, 6, 37, 55, 2]. A common practice in few-shot action recognition is to adopt a metric-based meta-learning strategy, where classification probabilities are computed based on the distance between query video features and support prototypes. Prior works [26, 4, 30, 43, 14, 20] achieve this by fine-tuning large image models, such as CLIP [31], on video datasets alongside independent temporal modules, as shown in Fig. 1 (a). Beyond visual-centric approaches, recent studies have explored multimodal frameworks for few-shot learning by integrating semantic information. For instance, works like [45, 17, 57] employ LLM or VLM to expand label descriptions, demonstrating broad utility.
Although previous studies have made progress, several limitations remain. First, full fine-tuning erodes the model’s inherent prior knowledge, thereby degrading its ability to generalize to unseen tasks. Second, the importance of incorporating task-specific adaptation into parameter-efficient fine-tuning (PEFT) is often overlooked, which leads to the neglect of the most discriminative information in each meta-task. Third, enhanced label information is typically treated as a static description, ignoring dynamic and ordered aspects that are essential for capturing semantic cues in fine-grained sequential modeling. Fourth, existing approaches inadequately align cross-modal features because they simply average features while disregarding temporal coupling relation. Motivated by these observations in FSAR, we propose Task-Adapter++, a parameter-efficient dual adaptation framework for both image and text encoder. This method comprehensively incorporates task-specific information for modeling video features and preserves the ordering information in enhanced action descriptions.
To address the drawback of full fine-tuning, we turn to the recently emerged PEFT approach. Originating from the field of NLP [13, 12, 16], PEFT aims to adapt large pre-trained models to downstream applications without fine-tuning all the parameters of the model, while attaining performance comparable to that of full fine-tuning. In the field of computer vision, a series of adapter-tuning works [54, 17] have demonstrated the effectiveness of PEFT for video understanding. Inspired by these works, we argue such methods are particularly well-suited for FSAR, since they enable better trade-offs between leveraging the generalizable prior knowledge of pre-trained models and incorporating domain-specific knowledge from downstream tasks.
On the basis of the existing PEFT, we hope to introduce task-specific adaptation into feature extraction. Therefore, we propose a novel adapter-based framework with task-specific adaptation capabilities. Unlike existing methods that apply task adaptation to extracted features, we uniquely integrate it into the feature extractor itself. As shown in Fig. 1 (b), by reusing the frozen self-attention blocks to perform cross-video task-specific attention, our approach dynamically highlights discriminative task-relevant patterns within the support set. This innovation bridges a critical gap in few-shot learning: balancing the preservation of generalizable pre-trained knowledge with the integration of domain-specific cues, enabling more robust and sample-efficient video understanding.
Thirdly, motivated by the observation that real-world events inherently follow stage-specific sub-actions, we leverage LLM to generate and structurally link three temporally ordered sub-actions (derived from input labels) that describe an action’s beginning, process and end. Unlike prior text-based enhancements that treat descriptions as unordered, our approach enforces sequential dependencies between sub-actions, aligning semantic hierarchies with the temporal structure of videos. Further, we propose semantic order adaptation, a novel method that explicitly models the hierarchical progression of actions through language, displayed in Fig. 1 (c). This enables the semantic branch to complement visual-temporal features with fine-grained, order-aware context—a critical advancement for bridging the modality gap in FSAR.
At last, we propose a fine-grained cross-modal alignment strategy tailored for temporally video features. Unlike prior methods that force global average feature, our method first finely models the adjacent frame and then matches cross-modal features by stages. To be specific, we enhance the similarity calculation that will be performed separately on the three semantic and visual features, shown in Fig 1 (c). This innovation addresses a key gap in FSAR: matching multimodal cues finely within structured temporal stages rather than treating actions as a whole unit.
In conclusion, overall of framework is presented in Section 3 and our contributions can be summarized into the following four points:
  • We introduce a novel parameter-efficient dual adaptation method called Task-Adapter++, which harnesses the unique advantages from both visual and semantic aspects with adapters, effectively alleviating the issues of catastrophic forgetting and overfitting induced by full fine-tuning.
  • To the best of our knowledge, we are the first to propose performing task-specific adaptation during the process of spatio-temporal feature extraction, which significantly enhances the discriminative features specific to the given few-shot learning task.
  • We extend the semantic information with dynamic sub-action description using LLM and further model the sequential relationship of actions by semantic order adaptation.
  • We propose a fine-grained cross-modal alignment method by modeling visual features, which helps mapping visual features to reside in the same temporal stage as semantic description.
  • We clearly improve the performance over the previous version, with an overall improvement up to 3.4% on 5 benchmarks. And additional experiments fully demonstrate the advanced nature and effectiveness of each component of our Task-adapter++.
In this paper, we extend our previously published conference paper in MM’24 [3] in the following main aspects. (i) We propose a dual adaptation method for few-shot action recognition, which fine-tune the adapters concurrently in both image and text encoders. (ii) Benefiting from the LLM, we extend the labels to more comprehensive and fine-grained sub-action descriptions. This step greatly enriches semantic information on the temporal level. (iii) We further propose an adaptation strategy for the semantic branch, namely semantic order adapter, which better models sequential relationships of sub-actions. (iv) We introduce a fine-grained cross-modal alignment strategy ensuring visual features resides in the same temporal stage with semantic features. (v) We demonstrate each component’s effectiveness by comprehensive ablation experimental analysis and the mechanism of module action is further elucidated. (vi) We improve the SOTA performance over 5 benchmarks.
Figure 2: AIM (a) adapts the standard ViT bock (b) by freezing the original pre-trained model (outlined with a blue background) and adding tunable Adapters (c) individually for temporal adaptation, spatial adaptation and joint adaptation.
Figure 3: Illustration of Task-Adapter++. Note that we separately add task adapters and semantic order adapters into the two branches, i.e., image encoder and text encoder. After feature extraction, support and query prototypes are passed to the metric module to compute the classification scores and the query prototypes are sent into a alignment module before the metric computation with semantic features. The final score can be calculated on these two.
2 Related Work
2.1 Few-shot Learning
Few-shot learning aims to learn a model that can quickly adapt to unseen classes using only a few labeled data. Existing works can be divided into three categories: augmentation-based [25, 46], metric-based [55, 2, 33, 39, 50, 7, 21, 62] and optimization-based methods [8, 53]. Augmentation-based methods use auxiliary data to enhance the feature representations or generate extra samples [25] to increase the diversity of the support set. Metric-based methods focus on learning a generalizable feature extractor and classify the query instance by measuring the feature distance. Optimization-based methods aim to learn a good network initialization which can fast adapt to unseen classes with minimal optimization steps. Among these approaches, metric-based methods are commonly adopted thanks to its simplicity and superior performance. Additionally, task-specific learning has been explored in image-domain metric-based works [32, 1, 48, 24, 18] to achieve better few-shot learning performance. FEAT [55] introduces set-to-set function to obtain task-specific visual features for the few-shot learning tasks. Ta-Adapter [59] introduces a novel approach that enhances few-shot image classification by incorporating task-aware CLIP encoders and a collaborative prompt learning strategy, demonstrating superior performance. However, these methods all introduce task-specific modules at the level of well-extracted image features, which is insufficient for more complex video features due to the additional temporal dimension. In this work, we follow the metric-based paradigm and perform a novel adaptation during the feature extraction process to achieve more comprehensive adaptation to video data.
2.2 Few-shot Action Recognition
In few-shot action recognition, most of the previous works [26, 58, 30, 66, 23, 68, 14, 35, 52, 27, 20] fall into the metric-based methods. To solve the problem of temporal misalignment when comparing different videos, a wide series of temporal alignment modules are proposed. OTAM [4] improves the dynamic time warping algorithm to preserve the temporal ordering inside the video feature similarities. TRX [30] proposes to take the ordered tuples of frames as the least matching unit instead of single frames. CMOT [23] uses the optimal transport algorithm to find the best matching between two video features. HyRSM [43] treats the sequence matching problem as a set matching problem for better alignment. Except for the aforementioned works, there are also studies focusing on spatial relations [60] and multi-modal fusion [9, 40, 47]. However, there is still a lack of research on task-specific learning in few-shot action recognition. In this work, we emphasize the vital importance of task-specific adaptation in identifying the most discriminative spatial-temporal features within the provided video samples.
2.3 Parameter-Efficient Fine-Tuning
Parameter-Efficient Fine-Tuning (PEFT) is derived from NLP [13, 12, 16], aiming to fine-tune large pre-trained models on downstream tasks in an efficient way while achieving comparable performance to full fine-tuning. CLIP-Adapter [10] adds a learnable bottleneck layer behind the frozen CLIP backbone to learn new features for downstream tasks. AIM [54] reuses the frozen pre-trained attention layer and applies it to the temporal dimension of the input video. SAFE [67] applies PEFT to semantic awareness that enhance the comprehension ability of image classification. In this work, we propose a novel PEFT method to acquire the most discriminative information and order information on two branches of CLIP.
3 Method
3.1 Problem definition
Few-shot action recognition aims to learn knowledge from base classes and classify human action on new classes with limited data. To achieve this goal, we adopt an episode-based training manner, which can be performed as $C$-way $K$-shot. In each episode, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from the training set $\mathcal{D}_{tr}$ for each episode during training. All the $C\times K$ video samples form the support set $\mathcal{S_{tr}}=\{x_{i}\}_{i=1}^{C\times K}$. Additionally, there is a query set $\mathcal{Q}_{tr}=\{x_{i}\}_{i=1}^{Q_{tr}}$ from the same $C$ action classes. The learning objective of each task is to classify each query video $q\in\mathcal{Q}_{tr}$ into one of the $C$ classes. In test stage, we sample $C$-way $K$-shot task from testing set $\mathcal{D}_{ts}$ and classify the query video $q\in\mathcal{Q}_{ts}$ into one of the $C$ classes with frozen parameters.
Figure 4: In task-specific adaptation, we introduce task adapter after T-MSA and S-MSA to enhance the task-specific information for the few-shot action recognition. This part illustrates the computational process of the image encoder given a 2-way 1-shot learning task with two query videos in the query set.
3.2 Framework
The architecture of whole Task-adapter++ is shown in Fig. 3, which reveal the main framework of our dual adaptation. The visual branch is shown in Fig. 4 where we introduce task specific adaptation to pay attention to the most discriminative video information. And Fig. 5 shows the semantic branch which model the sequential information between sub-actions by our proposed semantic order adaptation.
3.3 Adapting ViT to Action Recognition
Given an RGB video sample $v\in\mathbb{R}^{T\times H\times W\times 3}$, a vanilla ViT takes the temporal dimension $T$ as the batch dimension $B$, i.e., there is no exchange of information between frames. Specifically, ViT divides each frame into $N=HW/P^{2}$ patches, where $P$ is the patch size, and transforms the patches into $D$ dimensional patch embeddings $v_{p}\in\mathbb{R}^{T\times N\times D}$. A learnable [class] token is prepended to the patch embeddings $v_{p}$ as $v_{0}\in\mathbb{R}^{T\times(N+1)\times D}$. To retain position information, a learnable positional embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ is added to $v_{0}$ as $z_{0}=v_{0}+E_{pos}$, where the $E_{pos}$ is broadcast to $T$ dimension to be shared among different frames and $z_{0}$ is the final input of the stacked ViT blocks. Each ViT block consists of LayerNorm (LN), multi-head self-attention (MSA) and MLP layers, with residual connections after each block, as shown in Fig. 2 (a). The computation process during each ViT block can be given by:
$z_{l}^{\prime}$ $={\rm MSA}({\rm LN}(z_{l-1}))+z_{l-1}$ (1)
$z_{l}$ $={\rm MLP}({\rm LN}(z_{l}^{\prime}))+z_{l}^{\prime}$ (2)
where $z_{l}$ denotes the output of the $l$-th ViT block and all the MSAs are computed on patch dimension to learn the relationship of different spatial parts of the original frame. The [class] tokens of each frame from the last ViT block are concatenated as the final video feature $F_{v}\in\mathbb{R}^{T\times D}$ which is fed into the classifier and used to calculate the final classification score.
Inspired by PEFT techniques, AIM tries to adapt pre-trained image model to video action recognition with minimal fine-tuning by freezing the original model and only fine-tuning additionally introduced adapters shown in Fig. 2 (b). Moreover, AIM proposes to reuse the pre-trained MSA layer to do temporal modeling which is deficient in the vanilla ViT. As shown in Fig. 2 (c), a shared frozen temporal MSA (T-MSA) layer is prepended to the original spatial MSA (S-MSA) sequentially to exchange the temporal information between frames for temporal modeling, which can be achieved easily by permuting the patch dimension and the temporal dimension. The tunable Adapter is added after each MSA layer and parallel to the MLP layer to achieve temporal, spatial and joint adaptation as follows:
$z_{l}^{t}$ $={\rm Adapter}({\rm T\mbox{-}MSA}({\rm LN}(z_{l-1})))+z_{l-1}$ (3)
$z_{l}^{s}$ $={\rm Adapter}({\rm S\mbox{-}MSA}({\rm LN}(z_{l}^{t})))+z_{l}^{t}$ (4)
$z_{l}$ $={\rm Adapter}(({\rm LN}(z_{l}^{s}))+{\rm MLP}({\rm LN}(z_{l}^{s}))+z_{l}^{s}$ (5)
where $z_{l}^{t}$, $z_{l}^{s}$ and $z_{l}$ individually stands for temporal adapted, spatial adapted and jointly adapted output from the $l$-th ViT block. However, directly transferring this method to few-shot action recognition is inadequate due to the lack of considering the relationship between different videos within the few-shot learning task and the information gap between seen classes and unseen classes.
Figure 5: In semantic order adaptation, we introduce semantic order adapter after MSA to model the sequential information of sub-actions. This part illustrates the computational process of the text encoder given a 2-way 1-shot learning task.
3.4 Adapting ViT to Task-specific FSAR
FSAR studies how to classify unseen class query videos with only a limited number of labeled samples available. Following the episodic training, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from the training set $\mathcal{D}_{tr}$ for each episode during training. All the ${C\times K}$ video samples form the support set $\mathcal{S}=\{x_{i}\}_{i=1}^{CK}$. Additionally, there is a query set $\mathcal{Q}=\{x_{i}\}_{i=1}^{Q}$ from the same $C$ action classes. The learning objective of each task is to classify each query video $q\in\mathcal{Q}$ into one of the $C$ classes. Although the temporal adaptation proposed by AIM gives vanilla ViT the ability of temporal modeling, it remains unaddressed that the ViT model could not attend to different videos and enhance the task-specific discriminative features.
Next, we describe how we empower the frozen ViT with the ability to adapt embeddings with task-specific information for few-shot action recognition. As described in Section 3.3, we first map all the videos in the whole sampled task (both support set and query set) into their patch embeddings by Eq. (1) and Eq. (2), then we get the input features of the first ViT block $\mathcal{S}_{0}=[x_{0}^{1},...,x_{0}^{CK}]$ and $\mathcal{Q}_{0}=[x_{0}^{1},...,x_{0}^{Q}]$, where $x_{0}^{i}\in\mathbb{R}^{T\times(N+1)\times D}$ denotes the patch embeddings of the $i$-th video both for the support set and query set. Then, these patch embeddings are fed into several frozen ViT blocks to exchange spatial information between different tokens. It should be noted that we only introduce tunable adapters in the last L ViT blocks, which is a choice based on empirical study. We will further analyze the impact of this decision in the ablation experiment. Starting from the (N-L+1)-th layer, we introduce tunable adapter layers into the original model. Specifically, besides T-MSA and S-MSA, we further reuse the frozen MSA layer as Task-MSA after the temporal and spatial adaptation to perform task-specific self-attention across the tokens at the same spatial-temporal location in different videos (see Fig. 4). Take the support features as an example:
$\mathcal{S}_{l}^{t}$ $={\rm Adapter}({\rm T\mbox{-}MSA}({\rm LN}(\mathcal{S}_{l-1})))+\mathcal{S}_{l-1}$ (6)
$\mathcal{S}_{l}^{s}$ $={\rm Adapter}({\rm S\mbox{-}MSA}({\rm LN}(\mathcal{S}_{l}^{t})))+\mathcal{S}_{l}^{t}$ (7)
$\mathcal{S}_{l}^{task}$ $={\rm Adapter}({\rm Task\mbox{-}MSA}({\rm LN}(\mathcal{S}_{l}^{s})))+\mathcal{S}_{l}^{s}$ (8)
$\mathcal{S}_{l}$ $={\rm Adapter}({\rm LN}(\mathcal{S}_{l}^{task}))$
$+{\rm MLP}({\rm LN}(\mathcal{S}_{l}^{task}))+\mathcal{S}_{l}^{task}$ (9)
where $\mathcal{S}_{l-1},\mathcal{S}_{l}\in\mathbb{R}^{CK\times T\times(N+1)\times D}$ individually denotes the input and output support features of the $l$-th layer ($l\in\{N-L+1,N-L+2,...,N\}$).
The query features can be computed in the same way. To avoid exposing the target support class features to the specific query video feature, we isolate the information interaction between the support set and query set.
Finally, we concatenate the [class] tokens of each frame from the last ViT block to get the support set video representations $\mathcal{F}_{\mathcal{S}}\in\mathbb{R}^{CK\times T\times D}$ and the query set video representations $\mathcal{F}_{\mathcal{Q}}\in\mathbb{R}^{Q\times T\times D}$. Both the above prototypes will be used to calculate matching score in the metric step.
3.5 Semantic Order Adaptation
Essentially, our primary task is to construct a semantic corpus that contains dynamic descriptions of different actions. We call the LLM to generate fine-grained sub-actions by the temporal sequence with the following prompts:" Given an action label {LABEL}, you should describe three states of each action." For example, if the input label is long jump, the corresponding output is a list followed ["Run up for momentum.", "Take off with a powerful leap.", "Land and maintain balance."]
After acquiring the fine-grained descriptions above, we concatenate the original label with the three sub-actions to form three prompts, which are then used as input for the semantic branch of CLIP. Such as "A video of action about {label}: {sub-actions}".
Inspired by PEFT, we froze the text encoder and a tunable adapter is added after the last $M$-layer to achieve semantic order adaptation. We reuse the frozen MSA layers as O-MSA, which could model the 3 sub-actions. For instance:
$\mathcal{T}_{j^{\prime}}^{c}$ $={\mathrm{MSA}}({\mathrm{LN}}(\mathcal{T}_{j-1}^{c}))+\mathcal{T}_{j-1}^{c}$ (10)
$\mathcal{T}_{j^{\prime\prime}}^{c}$ $={\rm Adapter}({\rm O\mbox{-}MSA}({\rm LN}(\mathcal{T}_{j^{\prime}}^{c})))+\mathcal{T}_{j^{\prime}}^{c}$ (11)
$\mathcal{T}_{j}^{c}$ $={\rm MLP}({\rm LN}(\mathcal{T}_{j^{\prime\prime}}^{c}))+\mathcal{T}_{j^{\prime\prime}}^{c}$ (12)
where the output semantic feature of class $c$ in $j$-th ViT blockthe ($j\in\{N-M+1,N-M+2,...,N\}$) can be described as $\mathcal{T}_{j}^{c}=\{\mathcal{T}_{j}^{c_{1}},\mathcal{T}_{j}^{c_{2}},\mathcal{T}_{j}^{c_{3}}\}\in\mathbb{R}^{77\times 3\times D}$, where $77$ presents the context length in text encoder. After the whole branch, we concatenate the features of different actions within a task to obtain a vector $\mathcal{T}_{j}^{C}\in\mathbb{R}^{C\times 77\times 3\times D}$
Finally, we select the [class] token from semantic features $\mathcal{F}_{\mathcal{T}}\in\mathbb{R}^{C\times 3\times D}$ as the output followed by CLIP from the last layer of ViT block.
Figure 6: The structure of our alignment module. We use cross-attention as the alignment module. The ($i\mbox{-}1$)-th frame serves as $q$ and the $i$-th frame serves as $k$ and $v$.
3.6 Cross modal Alignment
Generally, considering that video has richer information and higher matching score than text, aligning video features to text is good choice. As shown in Fig. 6, we propose a novel alignment method that model the information between adjacent frames. The algorithm could be described as:
[…]
where $\mathcal{F}_{Q}^{i}\in\mathbb{R}^{C\times D}$ means the $i$-th frame of the input visual information, and for ease of understanding we define the aligned features as ¯$\bar{\mathcal{F}}_{Q}$.
The final question is how to calculate the classification result. The result is composed by two parts: video matching score of video prototypes and video-text matching score. For the support video feature $\mathcal{F}_{\mathcal{S}}$ and query video feature $\mathcal{F}_{\mathcal{Q}}$, the support video prototype can be denoted as:
^$\hat{\mathcal{F}}_{\mathcal{S}}^{c}=\frac{1}{K}\sum^{K}_{i}{\mathcal{F}_{\mathcal{S}}[c][i]}$ (14)
where ^$\hat{\mathcal{F}}_{\mathcal{S}}^{c}\in\mathbb{R}^{C\times T\times D}$ denotes the class prototype of the $c$-th support class. We directly average all the support video features as the class prototype for each class. To calculate the classification probability, we measure the distance between the query features and the class prototypes by an arbitrary metric module, e.g., TRX [30], OTAM [4] and Bi-MHM [43], which can formulated by:
$\mathcal{P}^{c}_{i}={\rm Metric}(\mathcal{F}_{\mathcal{Q}}[i],\hat{\mathcal{F}}_{\mathcal{S}})$ (15)
where $\mathcal{P}^{c}_{i}$ is the probability of the $c$-th class for the $i$-th query video in query set and ^$\hat{\mathcal{F}}_{\mathcal{S}}=[\hat{\mathcal{F}}_{\mathcal{S}}^{1},\hat{\mathcal{F}}_{\mathcal{S}}^{2},...,\hat{\mathcal{F}}_{\mathcal{S}}^{c}]\in\mathbb{R}^{C\times T\times D}$ stands for all the class prototypes of the whole support set. For the metric modules, we will demonstrate in our experiments that our proposed Task-Adapter++ consistently improves the performance of any metric methods. Besides that, owning to the superiority of our task-specific adaptation during feature extraction, excellent performance can be achieved simply by averaging the frame-to-frame cosine similarity without complex metric measurements.
We calculate the possibility distribution of cross-modality by cosine similarity using semantic features $\mathcal{F}_{\mathcal{T}}$ and aligned query features ¯$\bar{\mathcal{F}}_{\mathcal{Q}}$.
$\dot{\mathcal{P}}=\sum_{s=1}^{3}\left[\begin{array}[]{c}\mathcal{F}_{\mathcal{T}}^{1s}\ \mathcal{F}_{\mathcal{T}}^{2s}\ \cdots\ \mathcal{F}_{\mathcal{T}}^{Cs}\end{array}\right]\left[\begin{array}[]{c}\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s+1]}\\ \bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s+1]}\\ \vdots\\ \bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s+1]}\\ \end{array}\right]$ (16)
where $\mathcal{F}_{\mathcal{T}}^{Cs}$ means the $s$-th sub-action features of class $C$, and ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s]}$ means the $2s$-th frame aligned query feature in class $C$. The process is explained well by Fig. 7.
Figure 7: The illustration of matching process by stages. The aligned query features are divided into three segments and matched with the features of three sub-actions respectively.
Different from previous works, we cut query features into three pieces instead of simply averaging them. Therefore, the matching between query videos and sub-actions can be coupled in temporal dimension. We combine the zero-shot results and few-shot results in the same way as CLIP-FSAR [41] which uses a simple element-wise multiplication. Thus, the final score is calculated in the form of a product:
$\text{Score}=\mathcal{P}\times\dot{\mathcal{P}}/9$ (17)
During the training stage, classification probabilities for each task serve as the logits for computing cross-entropy loss and backpropagating gradients. Importantly, we solely fine-tune the parameters of the additional adapter layers while keeping the original pre-trained model frozen. In testing, we freeze the entire backbone, including the adapters, and directly extract features for the unseen test classes.
Table 1: This part reported the results focused on temporal-related dataset (SSv2-Small and SSv2-Full). Comparison with current SOTA few-shot action recognition methods on 5-way 1-shot, 5-way 3-shot and 5-way 5-shot benchmarks. The best results are highlighted in bold.
Method Reference Backbone Pretrain Fine-tuning SSv2-Small SSv2-Full
1-shot 3-shot 5-shot 1-shot 3-shot 5-shot
CMN [64] ECCV’18 ResNet50 IN-21K Full Fine-tuning 34.4 42.5 43.8 36.2 - 48.9
OTAM [4] CVPR’20 ResNet50 IN-21K Full Fine-tuning - - - 42.8 51.5 52.3
TRX [30] CVPR’21 ResNet50 IN-21K Full Fine-tuning 36.0 51.9 59.1 42.0 57.6 64.6
STRM [36] CVPR’22 ResNet50 IN-21K Full Fine-tuning 37.1 49.2 55.3 43.1 59.1 68.1
HyRSM [43] CVPR’22 ResNet50 IN-21K Full Fine-tuning 40.6 52.3 56.1 54.3 65.1 69.0
HCL [61] ECCV’22 ResNet50 IN-21K Full Fine-tuning 38.7 49.1 55.4 47.3 - -
MoLo [42] CVPR’23 ResNet50 IN-21K Full Fine-tuning 41.9 - 56.2 55.0 - 69.6
HyRSM++ [44] PR’24 ResNet50 IN-21K Full Fine-tuning 42.8 52.4 58.0 55.0 66.0 69.8
[…]
SA-CT [60] MM’23 ViT-B IN-21K Full Fine-tuning - - - - - 66.3
CLIP-FSAR [41] IJCV’23 ViT-B CLIP Full Fine-tuning 54.6 - 61.8 62.1 - 72.1
MA-CLIP [51] ArXiv’23 ViT-B CLIP PEFT 59.1 - 64.5 63.3 - 72.3
TSAM [17] Arxiv’24 ViT-B CLIP PRFT 60.5 - 66.7 65.8 - 74.6
Task-Adapter MM’24 ViT-B CLIP PEFT 60.2 66.1 70.2 71.3 71.5 74.2
Task-Adapter++ - ViT-B CLIP PEFT 63.6 69.1 73.2 72.8 73.0 75.5
4 Experiments
4.1 Experimental Settings
Datasets.We evaluate our method on four datasets that widely used in few-shot action recognition, namely HMDB51 [15], UCF101 [34], Kinetics [5] and SSv2 [11]. The first three datasets focus on scene understanding, while the last dataset, SSv2, has been shown to require more challenging temporal modeling [38, 4]. Particularly, We use the few-shot split proposed by [58] for HMDB51 and UCF101, with 31/10/10 classes and 70/10/21 classes for train/val/test, respectively. Following [64], Kinetics is used by selecting a subset which consists of 64, 12 and 24 training, validation and testing classes. For SSv2, we provide the results of two publicly available few-shot split, i.e., SSv2-Small [65] and SSv2-Full [4]. Both SSv2-Small and SSv2-Full contain 100 classes selected from the original dataset, with 64/12/24 classes for train/val/test, while SSv2-Full contains 10x more videos per class in the training set.
Implementation details. As previously illustrated, on one hand we aim to adapt the image pre-trained ViT model for task-specific few-shot action recognition, on the other hand we use order adaptation to model semantic information of sub-actions. To better demonstrate the model’s capabilities, we utilized a CLIP pre-trained ViT model [31] as the backbone. For a fair comparison with the existing works [41, 51], we uniformly sample 8 frames for each video, and crop the frames to 224$\times$224 as the input resolution. Before the training, we use GPT-4 [28] as LLM to enhance our label information. The semantic corpus is saved as json file without any manual modification. During the training stage, random crop is used for data augmentation. We freeze the original pre-trained model and only fine-tune the introduced adapters (as shown in Fig. 2 (c)) using the SGD optimizer with a learning rate of 0.001. For the testing stage, we freeze all the parameters of our backbone to extract task-specific discriminative video features and semantic features for measuring the distances. All reported results in our paper are the average accuracy over 10,000 few-shot action recognition tasks.
Table 2: This part reported the results focused on scene-related datasets (including HMDB51, UCF101, and Kinetics). Comparison with current SOTA few-shot action recognition methods on 5-way 1-shot, 5-way 3-shot and 5-way 5-shot benchmarks. The best results are highlighted in bold.
Method Reference Backbone Pretrain Fine-tuning HMDB51 UCF101 Kinetics
1-shot 3-shot 5-shot 1-shot 3-shot 5-shot 1-shot 3-shot 5-shot
CMN [64] ECCV’18 ResNet50 IN-21K Full Fine-tuning - - - - - - 60.5 75.6 78.9
ARN [58] ECCV’20 C3D - Full Fine-tuning 44.6 - 59.1 62.1 - 84.8 63.7 - 82.4
OTAM [4] CVPR’20 ResNet50 IN-21K Full Fine-tuning 54.5 65.7 68.0 79.9 87.0 88.9 73.0 78.7 85.8
TRX [30] CVPR’21 ResNet50 IN-21K Full Fine-tuning 53.1 67.4 75.6 78.2 92.4 96.1 63.6 81.8 85.9
STRM [36] CVPR’22 ResNet50 IN-21K Full Fine-tuning 52.3 67.4 77.3 80.5 92.7 96.9 62.9 81.1 86.7
HyRSM [43] CVPR’22 ResNet50 IN-21K Full Fine-tuning 60.3 71.7 76.0 83.9 93.0 94.7 73.7 83.5 86.1
HCL [61] ECCV’22 ResNet50 IN-21K Full Fine-tuning 59.1 71.2 76.3 82.5 91.0 93.9 73.7 82.4 85.8
MoLo [42] CVPR’23 ResNet50 IN-21K Full Fine-tuning 60.8 - 77.4 86.0 - 95.5 74.0 - 85.6
MGCSM [56] MM’23 ResNet50 IN-21K Full Fine-tuning 61.3 - 79.3 86.5 - 97.1 74.2 - 88.2
HyRSM++ [44] PR’24 ResNet50 IN-21K Full Fine-tuning 61.5 72.7 76.4 85.8 93.5 95.9 74.0 83.9 86.4
[…]
SA-CT [60] MM’23 ViT-B IN-21K Full Fine-tuning - - 81.6 - - 98.0 - - 91.2
CLIP-FSAR [41] IJCV’23 ViT-B CLIP Full Fine-tuning 77.1 - 87.7 97.0 - 99.1 94.8 - 95.4
MA-FSAR [51] ArXiv’23 ViT-B CLIP PEFT 83.4 - 87.9 97.2 - 99.2 95.7 - 96.0
TSAM [17] AAAI’25 ViT-B CLIP PEFT 84.5 - 88.9 98.3 - 99.3 96.2 - 97.1
Task-Adapter MM’24 ViT-B CLIP PEFT 83.6 88.0 88.8 98.0 98.7 99.4 95.0 96.3 96.8
Task-Adapter++ - ViT-B CLIP PEFT 86.5 90.6 92.0 98.9 99.5 99.7 96.0 97.3 97.5
4.2 Comparison with state-of-the-art methods
Table 1 and Table 2 presents a comprehensive comparison of our method with recent few-shot action recognition methods. As our Task-Adapter++ is designed for ViT-based architectures, we use CLIP [31] pre-trained ViT to make a fair comparison with recent few-shot action recognition works that use large-scale pretrained models. Of particular note, by comparing the results of works that use CLIP pre-trained models and those that use ImageNet (IN-21K) pre-trained models, we can conclude that the performance of few-shot action recognition indeed benefits from large-scale pretraining, especially for scene-related datasets.
As shown in Table 1, our Task-Adapter++ can achieve best results on SSv2-Small and SSv2-Full, indicating that our method can model the temporal information well. For Average 1-shot, 3-shot and 5-shot, our method outperforms the exsisting SOTA methods by 3.1% on SSv2-Small and 1.4% on SSv2-Full, respectively. Compared with full fine-tuning method CLIP-FSAR [41], we can see our method has an overall improvement (from 3.4% to 11.4%). Compared with PEFT method TSAM [17], our method also has an overrall improvement (from 0.9% to 7.0%) on two datasets. We generally believe that SSv2 requires more temporal modeling than other three datasets, making it much more challenging, and we have achieved a significant improvement of at least 3% on SSv2-Small compared to the SOTA, indicating that our task can better adapt to the classification of complex actions.
In addition, we display the results on HMDB51, UCF101 and Kinetics in Table 2. It is obvious that we achieve SOTA results than other methods. Compared with full fine-tuning method CLIP-FSAR [41], we achieve an average improvement of 6.9% on HMDB51, 1.3% on UCF101 and 1.7% on Kinetics. Compared with PEFT method TSAM [17], our method also has an average improvement of 2.6% on HMDB51, 0.5% on UCF101 and 0.1% on Kinetics. These scene-related datasets require strong background understand ability which means our model can handle complex scene knowledge well.
Table 3: We choose frozen backbone as the baseline and compare the performance of different settings for 5-way 1-shot task on SSv2-Small.
Frozen Backbone AIM Task-specific Adaptation Partial Adapting Acc
$\times$ $\times$ $\times$ 36.0
$\times$ $\times$ $\times$ $\times$ 47.8
$\times$ $\times$ 53.3
[…]
$\times$ 55.2
$\times$ 56.6
60.2
Table 4: We choose our Task-Adapter as the baseline and compare the performance of different settings for 5-way 1-shot task on SSv2-Small.
Task-Adapter Three Sub-actions Order Adaptation Cross Attention Alignment Acc
$\times$ $\times$ $\times$ 60.2
$\times$ $\times$ 62.0
$\times$ $\times$ 58.5
[…]
$\times$ 62.8
63.6
4.3 Ablation study
Effectiveness of each component. To demonstrate the effectiveness of each component, we perform detailed ablation experiments in Table 3 and Table 4. In Table 3, the first row shows the results of a frozen pre-trained model without any learnable adapters, while the second row refers to the performance of full fine-tuning. Comparing the results of the third and fourth rows, we can see the effectiveness of PEFT over traditional full fine-tuning in few-shot action recognition. Note that the "Partial Adapting" in the fourth column of Table 3 refers to a common strategy which only adds adapters in the last several ViT layers. We can observe from the table that the improvement of introducing task-specific adaptation into AIM is distinct (1.9%) when inserting adapters into all the layers of the pre-trained model. Additionally, using partial adapting strategy for AIM also brings an improvement of 1%. Even so, the improvement of introducing task-specific adaptation into AIM is notably enhanced under the partial adapting strategy, increasing from 1.9% to 5.9%. In Table 4, the first row presents the baseline result, while the second row shows the performance when applying the three sub-actions, which yields a 1.8% improvement. The third row reports the effect of applying cross attention alignment, leading to a slight decline of 1.7%. Notably, performing an alignment operation without enhancing semantic features is redundant, as the original Task-Adapter already accomplishes semantic alignment during feature extraction. This observation in the fourth row supports our assumption, where the model’s performance increases further from 1.8% to 2.3%. Incorporating order adaptation improves the performance of the Three Sub-actions. It models the sequential relationships between sub-actions in more detail. This change increases performance from 1.8% to 2.5% as shown in the fifth row. The final row summarizes the complete configuration.
Figure 8: Effect of inserting Task-Adapters into the last $L$ ViT layers (e.g., $L$ = 1, 2, 3, 6, 12) on scene-related datasets. The best is highlighted with a red star.
Figure 9: Effect of inserting semantic order adapters into the last $M$ ViT layers (e.g., $M$ = 1, 2, 3, …, 12) on 3 different datasets. The best is highlighted with a black star.
Table 5: The result that using different modalities of information.
methods visual only semantic only both
Task-Adapter 53.4 50.4 60.2
Task-Adapter++ 53.1 54.3 63.6
Figure 10: We demonstrate accuracy for each time step computed by a single sub-action description to prove its rationality.
Figure 11: The results of $N$-way 1-shot between different methods on SSv2-Small and Kinetics.
The effect of inserting position of adapters. We can see from the Fig. 8 that inserting our Task-Adapter into the top 2 layers is the best setting for all these three datasets. Since these scene-related datasets focus on background understanding and do not need much temporal modeling, a more lightweight setting with inserting adapters into only 2 top layers is more appropriate. In the Fig. 9, we demonstrate the impact of the insertion position of a semantic order adapter. Our analysis indicates that for datasets such as SSv2, which already encode rich action information within their original labels, employing a deep architecture for text sequence modeling is generally unnecessary. We posit that the labels in SSv2, characterized by an abundance of verbs and positional terms, combined with the enhanced contextual information from LLM, yield comprehensive representations. Conversely, for datasets with more straightforward action content, such as Kinetics and HMDB51, a more sophisticated modeling approach is warranted.
The result that using different modalities of information on SSv2-Small 5-way 1-shot. As shown in Table 5, we observe that Task-Adapter++ can significantly improve the classification accuracy of the semantic branch by 3.9%. This improvement is undoubtedly due to the proposed augmentations of semantic information. Our method effectively fills the gap of insufficient information in semantic visual matching, making the two legs of the model equally long.
The accuracy for each frames. In general, LLM provides three descriptions of sub-actions. Except for our claimed method in Section 3, it is easy to consider them as a whole. This actually includes two ideas: averaging three semantic features or concatenating the descriptions together as one prompt. Therefore, Fig. 10 shows the accuracy with different baselines, i.e., without using sub-actions description, averaging features and concatenating three descriptions together. Following Section 3, we separately compute accuracy for each temporal moment using the 3 sub-action descriptions. The results not only indicate that describe sub-actions in this manner is reasonable and helpful, but explain the information loss when just averaging or concatenating them.
$N$-way 1-shot results. Fig. 11 shows that our approach, compared to the original Task-Adapter, not only improves absolute accuracy but also better adapts to $N$-way 1-shot learning ($N$geq 5), i.e., experiences less relative decline. As $N$ increases, classification difficulty also increases, and model performance decline is inevitable. However, because our approach fully considers both semantic and visual features, the model will be less affected in SSv2-Small. And in Kinetics, the results are nearly saturated. It can be seen that the results descend slightly when $N$ increases, and the gap keeps very small.
Figure 12: The choice of different alignment modules on SSv2-Small 5-way 1-shot.
The exploration of different alignment modules. In Fig. 12, we explore different modules for the alignment task. We utilize multi-layer cross attention and 1D CNN (layer=1, stride=1, kernel=3, channel=512) to align visual features. The results indicate that stacking multi-layer cross attention is ineffective and does not improve results, while 1D CNN shows no advantages over cross attention.
Figure 13: Comparison of the performance achieved by combining different fine-tuning strategies with existing widely used metric measurement methods on SSv2-Small 5-way 1-shot.
Different fine-tuning strategies using different metric measurement methods. We also compare the performance of different fine-tuning strategies using different metric measurement methods to demonstrate the applicability of our Task-Adapter++. Four commonly used methods are taken in this paper, namely TRX [30], Bi-MHM [43], OTAM [4] and ProtoNet (e.g., directly averaging the frame-to-frame cosine similarity over temporal dimension without complex metric measurements). Taking full fine-tuning strategy as the baseline, we observe a consistent improvement across all metric measurement methods, as shown in Fig. 13. It should be noted that ProtoNet outperforms all other methods when uniformly using Task-Adapter++. We attribute this to the fact that with our Task-Adapter++, the model significantly improves the utilization rate of multimodal information and enhances fault tolerance on the basis of Task-Adapter.
Table 6: Effect of inserting position of task adapters and order adapters on SSv2-Small.
Task-Adapter Order Adapter
Position Tunable Param Acc Position Tunable Param Acc
All 14.1 M 55.2 All 1.58 M 61.1
Bottom 6 7.2 M 50.7 Bottom 6 790 K 62.1
Bottom 2 2.4 M xxx Bottom 2 263 K 62.6
Top 6 7.2 M 60.2 Top 6 790 K 61.2
Top 2 2.4 M 54.3 Top 2 263 K 63.6
Dual Partial Adapting. Firstly, we study the effect of different inserting positions of our Task-Adapter in Table 6 left. The default setting is inserting adapters into all of the 12 ViT layers. Alternatively, we also evaluate the settings which respectively insert adapters to the bottom 2/6 layers and top 2/6 layers of the model (assuming forward propagation is from bottom to top). Inserting adapters only into the top 6 layers has been shown to achieve comparable performance with inserting to all layers in AIM [54]. However, inserting our Task-Adapter only into the top 6 layers outperforms inserting it into all layers by a large margin (5%) for few-shot action recognition. We attribute this to the hypothesize that bottom layers focus on generic features that do not require extensive adaptation, while the top layers focus on task-specific and discriminative features that benefit a lot from our task-specific adaptation module. Additionally, We study the effect of different inserting positions of our semantic order adapter in Table 6 right. The default setting is inserting adapters into all of the 12 ViT layers. We evaluate settings inserting adapters to bottom 2/6 and top 2/6 layers. Notice that inserting adapters only top 2 layers achieves best performance with only 263 K parameters. Considering little difference in results at bottom 2/6 layers, we believe the semantic features have not been well extracted yet, which is exactly the opposite at the top 2/6 layers.
Table 7: Impact of position of Task-MSA and O-MSA inner one ViT layer on SSv2-Small 5-way 1-shot.
Task-MSA O-MSA
Position Acc Position Acc
in front of T&S-MSA 55.6 in front of MSA 62.2
between T&S-MSA 55.5 - -
back of T&S-MSA 60.2 back of MSA 63.6
Position of the Task-MSA and O-MSA inner ViT layer. In the visual branch, as illustrated in Section 3, the original AIM comprises two multi-head self-attention (MSA) modules, individually for spatial and temporal modeling, and a task adapter. We study the optimal placement of our proposed task-specific MSA (Task-MSA) relative to the existing two MSAs for improved task modeling. As shown in Table  7 left, inserting the Task-MSA after the temporal (T-) and spatial (S-) MSAs achieves the best performance. It is noteworthy that the Task-MSA consistently outperforms the baseline AIM [54], regardless of its insertion position. These results indicate that the introduction of the Task-MSA indeed enhances few-shot action recognition performance compared to AIM. Furthermore, placing the Task-MSA after the T- and S-MSAs enables maximum extraction of task-specific discriminative features. In the semantic branch, we retain the original Task-Adapter as our baseline without changes to the encoder. We find that relocating the order MSA (O-MSA) back of the original MSA yields higher scores than placing it frontally, as shown in Table 7 right. This is because we model the order relationship of sub-action sequences at the feature level, so positioning the O-MSA later enables better feature-level modeling. Nonetheless, both placements perform better than the baseline Task-Adapter, proving the effectiveness of the Order-MSA.
5 Conclusion
In this paper, we propose a novel method named Task-Adapter++ to better adapt the pre-trained image models for few-shot action recognition. Instead of using the traditional full fine-tuning strategy, we propose a parameter-efficient dual adaptation method that only fine-tune the newly inserted adapters which alleviates the issues of catastrophic forgetting and overfitting. Specifically, we design task-adapter module to enhance the most discriminative features within the given few-shot learning tasks in visual branch. We otherwise utilize semantic order adapter to model the sequence of sub-actions in semantic branch. We further apply cross attention for cross modal alignment operations, which remarkably improves the performance on five few-shot action recognition benchmarks and achieves the new state-of-the-art results, particularly on the challenging SSv2 dataset.
6 Acknowledgements
xxx

### v2

@@
-
+Task-Adapter++: Task-specific Adaptation with Order-aware Alignment for Few-shot Action Recognition
 Abstract
@@
 keywords:
+Few-shot action recognition, Parameter-efficient fine-tuning, Task-specific adaptation, Semantic order adaptation
+journal: Pattern Recognition
 […]
+[…]
+{graphicalabstract}
+{highlights}
+We propose Task-Adapter++, a parameter-efficient dual adaptation framework for few-shot action recognition, which alleviates catastrophic forgetting and overfitting induced by full fine-tuning.
+For the visual branch, we propose Task-Adapter to extract discriminative task-specific information during feature extraction.
+For the semantic branch, we propose semantic order adaptation, which utilizes dynamic descriptions to model the sequential relationships among sub-actions, thereby enhancing the performance of stage-wise cross-modal matching.
+Task-Adapter++ achieves state-of-the-art performance across five commonly-used few-shot action recognition benchmarks.
 Figure 1: Comparison of different paradigms.
 (a) Existing methods fully fine-tune the feature extractor and combine it with temporal alignment module.
-(b) We froze the image encoder and apply Task-Specific adaptation to extract the most discriminative information.
-(c) We use semactic order adapter to model sequential information within sub-actions and further calculate the score by stage with aligned visual features.
+(b) Existing methods extract semantic features from static enhanced descriptions and calculate the score by global matching.
+(c) For the image encoder, we freeze the backbone and apply task-specific adaptation to extract the most discriminative information.
+(d) For the text encoder, we use semantic order adapter to model sequential information within dynamic sub-actions and further calculate the score by stage with aligned visual features.
 1 Introduction
@@
 A common practice in few-shot action recognition is to adopt a metric-based meta-learning strategy, where classification probabilities are computed based on the distance between query video features and support prototypes.
-Prior works [26, 4, 30, 43, 14, 20] achieve this by fine-tuning large image models, such as CLIP [31], on video datasets alongside independent temporal modules, as shown in Fig.
+Prior works [14, 15, 16, 17] achieve this by fine-tuning large image models, such as CLIP [1], on video datasets alongside independent temporal modules, as shown in Fig.
 1 (a).
 Beyond visual-centric approaches, recent studies have explored multimodal frameworks for few-shot learning by integrating semantic information.
-For instance, works like [45, 17, 57] employ LLM or VLM to expand label descriptions, demonstrating broad utility.
+For instance, works like [18, 19] utilize the zero-shot classification ability of CLIP by video-text matching, significantly improving the accuracy of few-shot classification.
+Some works [20, 21, 22] additionally employ large language models (LLMs) or vision-language models (VLMs) to expand label information, enriching the attribute information of actions or objects, as shown in Fig.
+1 (b).
 Although previous studies have made progress, several limitations remain.
 First, full fine-tuning erodes the model’s inherent prior knowledge, thereby degrading its ability to generalize to unseen tasks.
-Second, the importance of incorporating task-specific adaptation into parameter-efficient fine-tuning (PEFT) is often overlooked, which leads to the neglect of the most discriminative information in each meta-task.
-Third, enhanced label information is typically treated as a static description, ignoring dynamic and ordered aspects that are essential for capturing semantic cues in fine-grained sequential modeling.
-Fourth, existing approaches inadequately align cross-modal features because they simply average features while disregarding temporal coupling relation.
-Motivated by these observations in FSAR, we propose Task-Adapter++, a parameter-efficient dual adaptation framework for both image and text encoder.
-This method comprehensively incorporates task-specific information for modeling video features and preserves the ordering information in enhanced action descriptions.
-To address the drawback of full fine-tuning, we turn to the recently emerged PEFT approach.
-Originating from the field of NLP [13, 12, 16], PEFT aims to adapt large pre-trained models to downstream applications without fine-tuning all the parameters of the model, while attaining performance comparable to that of full fine-tuning.
-In the field of computer vision, a series of adapter-tuning works [54, 17] have demonstrated the effectiveness of PEFT for video understanding.
-Inspired by these works, we argue such methods are particularly well-suited for FSAR, since they enable better trade-offs between leveraging the generalizable prior knowledge of pre-trained models and incorporating domain-specific knowledge from downstream tasks.
-On the basis of the existing PEFT, we hope to introduce task-specific adaptation into feature extraction.
-Therefore, we propose a novel adapter-based framework with task-specific adaptation capabilities.
-Unlike existing methods that apply task adaptation to extracted features, we uniquely integrate it into the feature extractor itself.
+Second, the importance of incorporating task-specific adaptation into spatio-temporal feature extraction is often overlooked, which leads to the neglect of the most discriminative information in each task.
+Third, the enhanced label information is typically treated as a static description, ignoring the different stages of action occurrence.
+Due to the absence of fine-grained temporal descriptions and semantic order modeling, it is difficult to fully exploit CLIP’s zero-shot classification capabilities in video understanding.
+Fourth, existing approaches inadequately align cross-modal features by global matching, which disregards the temporal coupling between visual and semantic features.
+Motivated by these observations, we introduce Task-Adapter++, a framework that optimally caters to few-shot action recognition (FSAR) through parameter-efficient dual adaptation.
+Our method comprehensively incorporates task-specific information during video feature extraction, and preserves semantic order information throughout both semantic modeling and cross-modal alignment.
+Specifically, to address the first drawback of full fine-tuning, we turn to the recently emerged parameter-efficient finetuning (PEFT) approach.
+Originating from the field of NLP, PEFT [23, 24, 25] aims to adapt large pre-trained models to downstream applications without fine-tuning all the parameters of the model, while attaining performance comparable to that of full fine-tuning.
+Inspired by these works, we argue such methods, which only fine-tune a small number of additional adapter parameters, are particularly well-suited for few-shot action recognition, since they balance the preservation of generalizable pre-trained knowledge with the integration of domain-specific cues, enabling more robust and data-efficient video understanding.
+As for the second drawback, existing adapter methods in video domain mainly focus on temporal modeling and lack the consideration of enhancing task-specific features for the few-shot learning tasks.
+Regarding human cognition, a common consensus is that the most discriminative features vary for different target tasks.
+For example, in distinguishing between “High Jump” and “Long Jump”, the body movements performed by the actor is more crucial, whereas in distinguishing between “High Jump” and “Pole Vault”, it is more significant whether the actor is holding a long pole.
+Extracting task-specific features has been a key issue in metric based few-shot image classification  [12, 26], while neglected in FSAR.
+Although a few works [16, 27] consider task-specific adaptation, they solely apply task-specific methods at the level of well-extracted features, which hinders the discovery of the most discriminative information for a given few-shot learning task.
+In this work, we propose a novel adapter-based method that brings the capability of task-specific adaptation directly into the process of video feature extraction.
 As shown in Fig.
-1 (b), by reusing the frozen self-attention blocks to perform cross-video task-specific attention, our approach dynamically highlights discriminative task-relevant patterns within the support set.
-This innovation bridges a critical gap in few-shot learning: balancing the preservation of generalizable pre-trained knowledge with the integration of domain-specific cues, enabling more robust and sample-efficient video understanding.
-Thirdly, motivated by the observation that real-world events inherently follow stage-specific sub-actions, we leverage LLM to generate and structurally link three temporally ordered sub-actions (derived from input labels) that describe an action’s beginning, process and end.
-Unlike prior text-based enhancements that treat descriptions as unordered, our approach enforces sequential dependencies between sub-actions, aligning semantic hierarchies with the temporal structure of videos.
-Further, we propose semantic order adaptation, a novel method that explicitly models the hierarchical progression of actions through language, displayed in Fig.
-1 (c).
-This enables the semantic branch to complement visual-temporal features with fine-grained, order-aware context—a critical advancement for bridging the modality gap in FSAR.
-At last, we propose a fine-grained cross-modal alignment strategy tailored for temporally video features.
-Unlike prior methods that force global average feature, our method first finely models the adjacent frame and then matches cross-modal features by stages.
-To be specific, we enhance the similarity calculation that will be performed separately on the three semantic and visual features, shown in Fig 1 (c).
-This innovation addresses a key gap in FSAR: matching multimodal cues finely within structured temporal stages rather than treating actions as a whole unit.
-In conclusion, overall of framework is presented in Section 3 and our contributions can be summarized into the following four points:
+1 (c), by reusing the frozen self-attention blocks with learnable adapters to perform cross-video task-specific attention, our approach dynamically highlights the most discriminative task-relevant patterns within each given task.
+Thirdly, motivated by the observation that real-world events inherently follow stage-specific sub-actions, we leverage LLMs to generate and structurally link temporally ordered sub-actions that describe an action’s beginning, process and end.
+Unlike prior text-based enhancements that regard descriptions as static constructs, our approach enforces sequential dependencies between sub-actions, aligning semantic sub-action with the temporal structure of videos.
+Furthermore, we propose semantic order adaptation, a novel method that explicitly models the sub-action processes underlying complex actions through language, as displayed in Fig.
+1 (d).
+This enables the semantic branch to provide fine-grained, order-aware context for the subsequent matching.
+Finally, when matching visual and semantic features, to overcome the limitations of global matching and fully take advantage of the temporal coupling between visual and semantic features, we propose a novel fine-grained cross-modal alignment strategy.
+This strategy aligns video features with semantic features across different stages.
+Specifically, after fine-grained temporal modeling, which is realized by an attention mechanism, we segment the visual features to directly correspond with the semantic features representing the beginning, process, and end stages.
+As shown in Fig.
+1 (d), the final cross-modal matching score is derived from this fine-grained, segment-wise alignment across different temporal stages.
+Figure 2: Illustration of Task-Adapter++.
+We first augment the labels with sub-action descriptions using LLMs, and then independently integrate task adapters and semantic order adapters into the visual and semantic branches.
+After feature extraction, the support and query features are input into a metric module to compute classification scores, while the query features are further processed through a cross-modal alignment module prior to cross-modal video-text matching.
+The final score is obtained by combining the outputs from video matching and video-text matching.
+As a whole, our framework consists of two branches and computes the matching score accordingly: a visual branch with an image encoder enhanced with task-specific adaptations, and a semantic branch with a text encoder incorporating semantic order adaptations.
+As shown in Fig.
+2, the visual matching score is computed based on the query video features and the support video prototypes extracted by the visual branch, while the cross-modal matching score is derived from the semantic features extracted by the semantic branch and the query video features after cross-modal alignment.
+The final score is obtained by multiplying these two parts.
+Our contributions can be summarized as follows:
+  • 1.
     We introduce a novel parameter-efficient dual adaptation method called Task-Adapter++, which harnesses the unique advantages from both visual and semantic aspects with adapters, effectively alleviating the issues of catastrophic forgetting and overfitting induced by full fine-tuning.
+  • 2.
     To the best of our knowledge, we are the first to propose performing task-specific adaptation during the process of spatio-temporal feature extraction, which significantly enhances the discriminative features specific to the given few-shot learning task.
-  • We extend the semantic information with dynamic sub-action description using LLM and further model the sequential relationship of actions by semantic order adaptation.
-  • We propose a fine-grained cross-modal alignment method by modeling visual features, which helps mapping visual features to reside in the same temporal stage as semantic description.
-  • We clearly improve the performance over the previous version, with an overall improvement up to 3.4% on 5 benchmarks.
-And additional experiments fully demonstrate the advanced nature and effectiveness of each component of our Task-adapter++.
-In this paper, we extend our previously published conference paper in MM’24 [3] in the following main aspects.
-(i) We propose a dual adaptation method for few-shot action recognition, which fine-tune the adapters concurrently in both image and text encoders.
-(ii) Benefiting from the LLM, we extend the labels to more comprehensive and fine-grained sub-action descriptions.
-This step greatly enriches semantic information on the temporal level.
-(iii) We further propose an adaptation strategy for the semantic branch, namely semantic order adapter, which better models sequential relationships of sub-actions.
-(iv) We introduce a fine-grained cross-modal alignment strategy ensuring visual features resides in the same temporal stage with semantic features.
-(v) We demonstrate each component’s effectiveness by comprehensive ablation experimental analysis and the mechanism of module action is further elucidated.
-(vi) We improve the SOTA performance over 5 benchmarks.
-Figure 2: AIM (a) adapts the standard ViT bock (b) by freezing the original pre-trained model (outlined with a blue background) and adding tunable Adapters (c) individually for temporal adaptation, spatial adaptation and joint adaptation.
-Figure 3: Illustration of Task-Adapter++.
-Note that we separately add task adapters and semantic order adapters into the two branches, i.e., image encoder and text encoder.
-After feature extraction, support and query prototypes are passed to the metric module to compute the classification scores and the query prototypes are sent into a alignment module before the metric computation with semantic features.
-The final score can be calculated on these two.
+  • 3.
+    We extend the semantic information with dynamic sub-action descriptions using LLMs and further model the sequential relationship of actions by semantic order adaptation.
+After that, we propose a fine-grained cross-modal alignment method which helps matching visual features with semantic features by stages.
+  • 4.
+    Extensive experiments fully demonstrate the superiority of our Task-Adapter++, establishing new state-of-the-art results on five standard few-shot action recognition benchmarks, and validating the effectiveness of each proposed component.
+In this paper, we extend our previously published conference paper in MM’24 [28] in the following aspects.
+(i) We propose a novel dual adaptation method for few-shot action recognition, which fine-tunes the adapters concurrently in both image and text encoders.
+(ii) Benefiting from the LLMs, we extend the labels to more comprehensive and fine-grained sub-action descriptions.
+This significantly enriches the semantic information at the temporal level.
+(iii) We further propose a novel adaptation strategy for the semantic branch, namely semantic order adapter, which better models the sequential relationships of sub-actions.
+(iv) We introduce a fine-grained cross-modal alignment strategy ensuring visual features reside in the same temporal stage as semantic features for better cross-modal feature matching.
+(v) We conduct comprehensive ablation studies to demonstrate the effectiveness of each proposed component of our Task-Adapter++.
+The underlying mechanism of each module is further elucidated, providing deeper insights into their contributions.
+(vi) We significantly boost the performance over the previous version, with an improvement up to 3.4% on 5 benchmarks.
 2 Related Work
@@
 Among these approaches, metric-based methods are commonly adopted thanks to its simplicity and superior performance.
-Additionally, task-specific learning has been explored in image-domain metric-based works [32, 1, 48, 24, 18] to achieve better few-shot learning performance.
+Additionally, task-specific learning has been explored in image-domain metric-based works [26, 39, 40, 41] to achieve better few-shot learning performance.
 FEAT [12] introduces set-to-set function to obtain task-specific visual features for the few-shot learning tasks.
@@
 2.2 Few-shot Action Recognition
-In few-shot action recognition, most of the previous works [26, 58, 30, 66, 23, 68, 14, 35, 52, 27, 20] fall into the metric-based methods.
+In few-shot action recognition, most of the previous works [15, 17, 43, 44, 45, 46, 47, 48] fall into the metric-based methods.
 To solve the problem of temporal misalignment when comparing different videos, a wide series of temporal alignment modules are proposed.
@@
 However, there is still a lack of research on task-specific learning in few-shot action recognition.
-In this work, we emphasize the vital importance of task-specific adaptation in identifying the most discriminative spatial-temporal features within the provided video samples.
+In this work, we emphasize the vital importance of task-specific adaptation in identifying the most discriminative spatio-temporal features within the provided video samples.
 2.3 Parameter-Efficient Fine-Tuning
@@
 SAFE [53] applies PEFT to semantic awareness that enhance the comprehension ability of image classification.
-In this work, we propose a novel PEFT method to acquire the most discriminative information and order information on two branches of CLIP.
+However, in the FSAR task, no work has ever considered extracting task-specific visual features and modeling the sequential information of sub-actions using PEFT.
+In this work, we propose a novel PEFT method to acquire the most discriminative information and order information on two branches of CLIP for FSAR.
 3 Method
@@
 To achieve this goal, we adopt an episode-based training manner, which can be performed as $C$-way $K$-shot.
-In each episode, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from the training set $\mathcal{D}_{tr}$ for each episode during training.
+The video datasets are partitioned into nonoverlapping training and testing splits, denoted as $\mathcal{D}_{tr}$ and $\mathcal{D}_{ts}$, respectively.
+In each episode, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from $\mathcal{D}_{tr}$ for each episode during training.
 All the $C\times K$ video samples form the support set $\mathcal{S}_{tr}=\{x_{i}\}_{i=1}^{C\times K}$.
@@
 The learning objective of each task is to classify each query video $q\in\mathcal{Q}_{tr}$ into one of the $C$ classes.
-In test stage, we sample $C$-way $K$-shot task from testing set $\mathcal{D}_{ts}$ and classify the query video $q\in\mathcal{Q}_{ts}$ into one of the $C$ classes with frozen parameters.
-Figure 4: In task-specific adaptation, we introduce task adapter after T-MSA and S-MSA to enhance the task-specific information for the few-shot action recognition.
+In test stage, we sample $C$-way $K$-shot task from $\mathcal{D}_{ts}$ and classify the query video $q\in\mathcal{Q}_{ts}$ into one of the $C$ classes with frozen parameters.
+Figure 3: Illustration of Task-Adapter in visual branch.
+For task-specific adaptation, we introduce task adapter after T-MSA and S-MSA to enhance the task-specific information for the few-shot action recognition.
 This part illustrates the computational process of the image encoder given a 2-way 1-shot learning task with two query videos in the query set.
 3.2 Framework
-The architecture of whole Task-adapter++ is shown in Fig.
-3, which reveal the main framework of our dual adaptation.
+The architecture of the whole Task-Adapter++ is shown in Fig.
+2, which demonstrate our dual adaptation method.
 The visual branch is shown in Fig.
-4 where we introduce task specific adaptation to pay attention to the most discriminative video information.
+3 where we introduce task-specific adaptation to pay attention to the most discriminative video information in a given task.
 And Fig.
 5 shows the semantic branch which model the sequential information between sub-actions by our proposed semantic order adaptation.
+The classification score consists of two parts: a metric score derived from the support and query features, and a cross-modal matching score computed from the query and semantic features.
+The details of each component are described as follows.
+Figure 4: AIM (a) adapts the standard ViT bock (b) by freezing the original pre-trained model (outlined with a blue background) and adding tunable Adapters (c) individually for temporal adaptation, spatial adaptation and joint adaptation.
 3.3 Adapting ViT to Action Recognition
@@
 Each ViT block consists of LayerNorm (LN), multi-head self-attention (MSA) and MLP layers, with residual connections after each block, as shown in Fig.
-2 (a).
+4 (b).
 The computation process during each ViT block can be given by:
@@
 Inspired by PEFT techniques, AIM tries to adapt pre-trained image model to video action recognition with minimal fine-tuning by freezing the original model and only fine-tuning additionally introduced adapters shown in Fig.
-2 (b).
+4 (c).
 Moreover, AIM proposes to reuse the pre-trained MSA layer to do temporal modeling which is deficient in the vanilla ViT.
 As shown in Fig.
-2 (c), a shared frozen temporal MSA (T-MSA) layer is prepended to the original spatial MSA (S-MSA) sequentially to exchange the temporal information between frames for temporal modeling, which can be achieved easily by permuting the patch dimension and the temporal dimension.
+4 (a), a shared frozen temporal MSA (T-MSA) layer is prepended to the original spatial MSA (S-MSA) sequentially to exchange the temporal information between frames for temporal modeling, which can be achieved easily by permuting the patch dimension and the temporal dimension.
 The tunable adapter is added after each MSA layer and parallel to the MLP layer to achieve temporal, spatial and joint adaptation as follows:
@@
 However, directly transferring this method to few-shot action recognition is inadequate due to the lack of considering the relationship between different videos within the few-shot learning task and the information gap between seen classes and unseen classes.
-Figure 5: In semantic order adaptation, we introduce semantic order adapter after MSA to model the sequential information of sub-actions.
-This part illustrates the computational process of the text encoder given a 2-way 1-shot learning task.
 3.4 Adapting ViT to Task-specific FSAR
 FSAR studies how to classify unseen class query videos with only a limited number of labeled samples available.
-Following the episodic training, we randomly sample a $C$-way $K$-shot task, which has $C$ action classes and $K$ labeled samples in each class, from the training set $\mathcal{D}_{tr}$ for each episode during training.
-All the ${C\times K}$ video samples form the support set $\mathcal{S}=\{x_{i}\}_{i=1}^{CK}$.
-Additionally, there is a query set $\mathcal{Q}=\{x_{i}\}_{i=1}^{Q}$ from the same $C$ action classes.
-The learning objective of each task is to classify each query video $q\in\mathcal{Q}$ into one of the $C$ classes.
 Although the temporal adaptation proposed by AIM gives vanilla ViT the ability of temporal modeling, it remains unaddressed that the ViT model could not attend to different videos and enhance the task-specific discriminative features.
-Next, we describe how we empower the frozen ViT with the ability to adapt embeddings with task-specific information for few-shot action recognition.
+Therefore, we describe how we empower the frozen ViT with the ability to adapt embeddings with task-specific information for few-shot action recognition.
 As described in Section 3.3, we first map all the videos in the whole sampled task (both support set and query set) into their patch embeddings by Eq.
@@
 Starting from the (N-L+1)-th layer, we introduce tunable adapter layers into the original model.
-Specifically, besides T-MSA and S-MSA, we further reuse the frozen MSA layer as Task-MSA after the temporal and spatial adaptation to perform task-specific self-attention across the tokens at the same spatial-temporal location in different videos (see Fig.
-4).
+Specifically, besides T-MSA and S-MSA, we further reuse the frozen MSA layer as Task-MSA after the temporal and spatial adaptation to perform task-specific self-attention across the tokens at the same spatio-temporal location in different videos (see Fig.
+3).
 Take the support features as an example:
@@
 $\mathcal{S}_{l}^{task}$ $={\rm Adapter}({\rm Task\mbox{-}MSA}({\rm LN}(\mathcal{S}_{l}^{s})))+\mathcal{S}_{l}^{s}$ (8)
-$\mathcal{S}_{l}$ $={\rm Adapter}({\rm LN}(\mathcal{S}_{l}^{task}))$
-$+{\rm MLP}({\rm LN}(\mathcal{S}_{l}^{task}))+\mathcal{S}_{l}^{task}$ (9)
+$\mathcal{S}_{l}$ $={\rm Adapter}({\rm LN}(\mathcal{S}_{l}^{task}))+{\rm MLP}({\rm LN}(\mathcal{S}_{l}^{task}))+\mathcal{S}_{l}^{task}$ (9)
 where $\mathcal{S}_{l-1},\mathcal{S}_{l}\in\mathbb{R}^{CK\times T\times(N+1)\times D}$ individually denotes the input and output support features of the $l$-th layer ($l\in\{N-L+1,N-L+2,...,N\}$).
@@
 Finally, we concatenate the [class] tokens of each frame from the last ViT block to get the support set video representations $\mathcal{F}_{\mathcal{S}}\in\mathbb{R}^{CK\times T\times D}$ and the query set video representations $\mathcal{F}_{\mathcal{Q}}\in\mathbb{R}^{Q\times T\times D}$.
-Both the above prototypes will be used to calculate matching score in the metric step.
-3.5 Semantic Order Adaptation
-Essentially, our primary task is to construct a semantic corpus that contains dynamic descriptions of different actions.
-We call the LLM to generate fine-grained sub-actions by the temporal sequence with the following prompts:" Given an action label {LABEL}, you should describe three states of each action." For example, if the input label is long jump, the corresponding output is a list followed ["Run up for momentum.", "Take off with a powerful leap.", "Land and maintain balance."]
-After acquiring the fine-grained descriptions above, we concatenate the original label with the three sub-actions to form three prompts, which are then used as input for the semantic branch of CLIP.
-Such as "A video of action about {label}: {sub-actions}".
-Inspired by PEFT, we froze the text encoder and a tunable adapter is added after the last $M$-layer to achieve semantic order adaptation.
-We reuse the frozen MSA layers as O-MSA, which could model the 3 sub-actions.
-For instance:
-$\mathcal{T}_{j^{\prime}}^{c}$ $={\mathrm{MSA}}({\mathrm{LN}}(\mathcal{T}_{j-1}^{c}))+\mathcal{T}_{j-1}^{c}$ (10)
-$\mathcal{T}_{j^{\prime\prime}}^{c}$ $={\rm Adapter}({\rm O\mbox{-}MSA}({\rm LN}(\mathcal{T}_{j^{\prime}}^{c})))+\mathcal{T}_{j^{\prime}}^{c}$ (11)
-$\mathcal{T}_{j}^{c}$ $={\rm MLP}({\rm LN}(\mathcal{T}_{j^{\prime\prime}}^{c}))+\mathcal{T}_{j^{\prime\prime}}^{c}$ (12)
-where the output semantic feature of class $c$ in $j$-th ViT blockthe ($j\in\{N-M+1,N-M+2,...,N\}$) can be described as $\mathcal{T}_{j}^{c}=\{\mathcal{T}_{j}^{c_{1}},\mathcal{T}_{j}^{c_{2}},\mathcal{T}_{j}^{c_{3}}\}\in\mathbb{R}^{77\times 3\times D}$, where $77$ presents the context length in text encoder.
-After the whole branch, we concatenate the features of different actions within a task to obtain a vector $\mathcal{T}_{j}^{C}\in\mathbb{R}^{C\times 77\times 3\times D}$
-Finally, we select the [class] token from semantic features $\mathcal{F}_{\mathcal{T}}\in\mathbb{R}^{C\times 3\times D}$ as the output followed by CLIP from the last layer of ViT block.
-Figure 6: The structure of our alignment module. We use cross-attention as the alignment module. The ($i\mbox{-}1$)-th frame serves as $q$ and the $i$-th frame serves as $k$ and $v$.
-3.6 Cross modal Alignment
-Generally, considering that video has richer information and higher matching score than text, aligning video features to text is good choice.
-As shown in Fig.
-6, we propose a novel alignment method that model the information between adjacent frames.
-The algorithm could be described as:
-[…]
-where $\mathcal{F}_{Q}^{i}\in\mathbb{R}^{C\times D}$ means the $i$-th frame of the input visual information, and for ease of understanding we define the aligned features as ¯$\bar{\mathcal{F}}_{Q}$.
-The final question is how to calculate the classification result.
-The result is composed by two parts: video matching score of video prototypes and video-text matching score.
-For the support video feature $\mathcal{F}_{\mathcal{S}}$ and query video feature $\mathcal{F}_{\mathcal{Q}}$, the support video prototype can be denoted as:
+The video metric score is calculated based on the support prototypes and query features.
+For the support video features $\mathcal{F}_{\mathcal{S}}$, the support video prototype can be denoted as:
 ^$\hat{\mathcal{F}}_{\mathcal{S}}^{c}=\frac{1}{K}\sum^{K}_{i}{\mathcal{F}_{\mathcal{S}}[c][i]}$ (10)
-where ^$\hat{\mathcal{F}}_{\mathcal{S}}^{c}\in\mathbb{R}^{C\times T\times D}$ denotes the class prototype of the $c$-th support class.
+where ^$\hat{\mathcal{F}}_{\mathcal{S}}^{c}\in\mathbb{R}^{T\times D}$ denotes the class prototype of the $c$-th support class.
 We directly average all the support video features as the class prototype for each class.
@@
 where $\mathcal{P}^{c}_{i}$ is the probability of the $c$-th class for the $i$-th query video in query set and ^$\hat{\mathcal{F}}_{\mathcal{S}}=[\hat{\mathcal{F}}_{\mathcal{S}}^{1},\hat{\mathcal{F}}_{\mathcal{S}}^{2},...,\hat{\mathcal{F}}_{\mathcal{S}}^{c}]\in\mathbb{R}^{C\times T\times D}$ stands for all the class prototypes of the whole support set.
-For the metric modules, we will demonstrate in our experiments that our proposed Task-Adapter++ consistently improves the performance of any metric methods.
+In our experiments of different metric modules, we will demonstrate that our proposed Task-Adapter++ consistently improves the performance of any metric methods.
 Besides that, owning to the superiority of our task-specific adaptation during feature extraction, excellent performance can be achieved simply by averaging the frame-to-frame cosine similarity without complex metric measurements.
-We calculate the possibility distribution of cross-modality by cosine similarity using semantic features $\mathcal{F}_{\mathcal{T}}$ and aligned query features ¯$\bar{\mathcal{F}}_{\mathcal{Q}}$.
-$\dot{\mathcal{P}}=\sum_{s=1}^{3}\left[\begin{array}[]{c}\mathcal{F}_{\mathcal{T}}^{1s}\ \mathcal{F}_{\mathcal{T}}^{2s}\ \cdots\ \mathcal{F}_{\mathcal{T}}^{Cs}\end{array}\right]\left[\begin{array}[]{c}\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{1}}^{[2s+1]}\\ \bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{2}}^{[2s+1]}\\ \vdots\\ \bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s-1]}+\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s]}+\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s+1]}\\ \end{array}\right]$ (16)
-where $\mathcal{F}_{\mathcal{T}}^{Cs}$ means the $s$-th sub-action features of class $C$, and ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{C}}^{[2s]}$ means the $2s$-th frame aligned query feature in class $C$.
-The process is explained well by Fig.
+Figure 5: Illustration of order adaptation in semantic branch.
+For semantic order adaptation, we introduce semantic order adapter after MSA to model the sequential information of sub-actions.
+This part illustrates the computational process of the text encoder given a 2-way 1-shot learning task.
+3.5 Semantic Order Adaptation
+Considering that existing works overlook the fine-grained sequential modeling in the semantic feature extraction process, we propose the semantic order adaptation.
+Essentially, our primary task is to construct a semantic corpus that contains dynamic descriptions of different actions.
+We prompt the LLMs to generate fine-grained sub-actions with following instruction: “Given an action label {LABEL}, describe three stages of the action.” The LLMs generate corresponding descriptions of the beginning, process, and end for each input action label.
+For example, if the input label is “Long Jump”, the corresponding output is a list delineates the beginning, process, and end of the action as follows: [“Run up for momentum.”, “Take off with a powerful leap.”, “Land and maintain balance.”]
+After obtaining the fine-grained descriptions as above, we concatenate the original label with the three sub-actions respectively to form three prompts, which are then used as input for the semantic branch of CLIP, such as “A video of action about {LABEL}: {SUB-ACTIONS-1}”, “A video of action about {LABEL}: {SUB-ACTIONS-2}” and “A video of action about {LABEL}: {SUB-ACTIONS-3}”.
+As shown in Fig.
+5, to enhance the representation of sequential relationships between sub-actions, inspired by PEFT, we freeze the text encoder and introduce a tunable adapter after the last $M$-layer to achieve semantic order adaptation.
+Specifically, we reuse the frozen MSA layers as O-MSA, which could model the sequential information across the three sub-actions as follows:
+$\mathcal{T}_{j^{\prime}}^{c}$ $={\mathrm{MSA}}({\mathrm{LN}}(\mathcal{T}_{j-1}^{c}))+\mathcal{T}_{j-1}^{c}$ (12)
+$\mathcal{T}_{j^{\prime\prime}}^{c}$ $={\rm Adapter}({\rm O\mbox{-}MSA}({\rm LN}(\mathcal{T}_{j^{\prime}}^{c})))+\mathcal{T}_{j^{\prime}}^{c}$ (13)
+$\mathcal{T}_{j}^{c}$ $={\rm MLP}({\rm LN}(\mathcal{T}_{j^{\prime\prime}}^{c}))+\mathcal{T}_{j^{\prime\prime}}^{c}$ (14)
+where the output semantic feature of class $c$ in the $j$-th ViT block ($j\in\{N-M+1,N-M+2,...,N\}$) is denoted as $\mathcal{T}_{j}^{c}=\{\mathcal{T}_{j}^{c_{1}},\mathcal{T}_{j}^{c_{2}},\mathcal{T}_{j}^{c_{3}}\}\in\mathbb{R}^{l\times 3\times D}$, and $l$ presents the sequence length in text encoder.
+Finally, for each action within a task, we select the end-of-text token from the text features in the last block and concatenate them to form the output semantic feature for cross-modal matching.
+Figure 6: The structure of our fine-grained cross-modal alignment.
+We use cross-attention for fine-grained temporal modeling, where the ($i\mbox{-}1$)-th frame serves as $q$ and the $i$-th frame serves as $k$ and $v$.
+3.6 Fine-grained Cross-modal Alignment
+Considering that videos contain richer spatio-temporal information while semantics are more abstract, mapping videos into the semantic space aligns more naturally with the principles of information extraction.
+Thus, we propose a fine-grained cross-modal alignment framework that models the temporal relationships between adjacent frames.
+In this framework, visual features are aligned to correspond with the same temporal stages as their semantic descriptions.
+As illustrated in Fig.
+6, our method enables more precise alignment by modeling frame-to-frame dynamics in the temporal context with cross-attention:
+[…]
+where $\mathcal{F}_{Q}^{i}\in\mathbb{R}^{Q\times D}$ denotes the $i$-th frame of the query’s visual feature set, and for ease of understanding we define the aligned features as ¯$\bar{\mathcal{F}}_{Q}\in\mathbb{R}^{Q\times(T-1)\times D}$.
+By default, $T$ is set to 8.
+As no padding is applied along the temporal dimension, the length of the aligned features is reduced by one compared to the original input sequence..
+For the $q$-th aligned visual features ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}$ in the query set, we divide it into three temporal segments, corresponding to the three different sub-actions, referred to as:
+¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}=[\frac{1}{3}\sum^{3}_{i=1}\bar{\mathcal{F}}_{\mathcal{Q}_{q}}^{i},\frac{1}{3}\sum^{5}_{i=3}\bar{\mathcal{F}}_{\mathcal{Q}_{q}}^{i},\frac{1}{3}\sum^{7}_{i=5}\bar{\mathcal{F}}_{\mathcal{Q}_{q}}^{i}]\in\mathbb{R}^{3\times D}$ (16)
+Then, the probability distribution is calculated based on the cosine similarity between the semantic feature $\mathcal{F}_{\mathcal{T}}$ and the aligned query feature segment ¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}$, which can be formulated as:
+$\dot{\mathcal{P}}_{q}^{c}=\frac{1}{3}\sum_{s=1}^{3}\text{Cos}(\bar{\mathcal{F}}_{\mathcal{Q}_{q}}[s],\mathcal{F}_{\mathcal{T}}^{c}[s])$ (17)
+where $\dot{\mathcal{P}}_{q}^{c}$ is the probability of the $c$-th class for the $q$-th video in query set.
+¯$\bar{\mathcal{F}}_{\mathcal{Q}_{q}}[s]$ means the $s$-th query feature segment for the $q$-th query video and $\mathcal{F}_{\mathcal{T}}^{c}[s]$ means the $s$-th sub-action features of class $c$.
+The process is also illustrated in Fig.
 7.
-Figure 7: The illustration of matching process by stages. The aligned query features are divided into three segments and matched with the features of three sub-actions respectively.
-Different from previous works, we cut query features into three pieces instead of simply averaging them.
-Therefore, the matching between query videos and sub-actions can be coupled in temporal dimension.
-We combine the zero-shot results and few-shot results in the same way as CLIP-FSAR [41] which uses a simple element-wise multiplication.
+Figure 7: The stage-wise illustration of the cross-modal feature matching process.
+The aligned query features are divided into three segments and matched with the features of three sub-actions respectively.
+Notably, instead of averaging temporal video features for global matching, we divide them into three temporal segments to facilitate fine-grained cross-modal alignment.
+This design enables the visual and textual features to be matched at each corresponding stage individually, resulting in more precise and discriminative matching result.
+3.7 Calculation of Classification Results
+The final classification score consists of two components: the video matching score $\mathcal{P}$, as demonstrated in Section 3.4, which measures the similarity between support and query features based on video prototypes; and the video-text matching score $\mathcal{\dot{P}}$, introduced in Section 3.6, which evaluates the alignment between semantic features and query features.
+These two scores are combined via element-wise multiplication to jointly consider both visual and semantic similarities.
 Thus, the final score is calculated in the form of a product:
-$\text{Score}=\mathcal{P}\times\dot{\mathcal{P}}/9$ (17)
-During the training stage, classification probabilities for each task serve as the logits for computing cross-entropy loss and backpropagating gradients.
-Importantly, we solely fine-tune the parameters of the additional adapter layers while keeping the original pre-trained model frozen.
-In testing, we freeze the entire backbone, including the adapters, and directly extract features for the unseen test classes.
-Table 1: This part reported the results focused on temporal-related dataset (SSv2-Small and SSv2-Full).
-Comparison with current SOTA few-shot action recognition methods on 5-way 1-shot, 5-way 3-shot and 5-way 5-shot benchmarks.
+$\text{Score}=\mathcal{P}\times\dot{\mathcal{P}}$ (18)
+The similarity scores for each task are treated as logits and used to compute the cross-entropy loss, enabling gradient backpropagation during training.
+Notably, we only fine-tune the parameters of the additional adapter layers, while keeping the backbone of the pre-trained model frozen.
+During testing, the entire model, including the adapters, is fixed, and the highest-scoring class for each query video is selected as the final prediction.
+Table 1: Comparison with current SOTA few-shot action recognition methods on temporal-related datasets (SSv2-Small and SSv2-Full) with 5-way 1-shot, 5-way 3-shot and 5-way 5-shot settings.
 The best results are highlighted in bold.
@@
 HyRSM++ [58] PR’24 ResNet50 IN-21K Full Fine-tuning 42.8 52.4 58.0 55.0 66.0 69.8
-[…]
+MASTAF [59] WACV’23 ViViT JFT Full Fine-tuning 45.6 - - 60.7 - -
 SA-CT [49] MM’23 ViT-B IN-21K Full Fine-tuning - - - - - 66.3
 CLIP-FSAR [60] IJCV’23 ViT-B CLIP Full Fine-tuning 54.6 - 61.8 62.1 - 72.1
-MA-CLIP [51] ArXiv’23 ViT-B CLIP PEFT 59.1 - 64.5 63.3 - 72.3
-TSAM [17] Arxiv’24 ViT-B CLIP PRFT 60.5 - 66.7 65.8 - 74.6
+MA-FSAR [19] ArXiv’23 ViT-B CLIP PEFT 59.1 - 64.5 63.3 - 72.3
+TSAM [21] AAAI’25 ViT-B CLIP PEFT 60.5 - 66.7 65.8 - 74.6
 Task-Adapter MM’24 ViT-B CLIP PEFT 60.2 66.1 70.2 71.3 71.5 74.2
@@
 4.1 Experimental Settings
-Datasets.We evaluate our method on four datasets that widely used in few-shot action recognition, namely HMDB51 [15], UCF101 [34], Kinetics [5] and SSv2 [11].
+Datasets.
+We evaluate our method on five datasets that widely used in few-shot action recognition, namely HMDB51  [61], UCF101 [62], Kinetics [7] and SSv2 [8].
 The first three datasets focus on scene understanding, while the last dataset, SSv2, has been shown to require more challenging temporal modeling [63, 14].
@@
 As previously illustrated, on one hand we aim to adapt the image pre-trained ViT model for task-specific few-shot action recognition, on the other hand we use order adaptation to model semantic information of sub-actions.
-To better demonstrate the model’s capabilities, we utilized a CLIP pre-trained ViT model [31] as the backbone.
+To better demonstrate the model’s capabilities, we utilize a CLIP pre-trained ViT model [1] as the backbone.
 For a fair comparison with the existing works [60, 19], we uniformly sample 8 frames for each video, and crop the frames to 224$\times$224 as the input resolution.
-Before the training, we use GPT-4 [28] as LLM to enhance our label information.
+Before the training, we use GPT-4 [65] to enhance our label information.
 The semantic corpus is saved as json file without any manual modification.
 During the training stage, random crop is used for data augmentation.
-We freeze the original pre-trained model and only fine-tune the introduced adapters (as shown in Fig.
-2 (c)) using the SGD optimizer with a learning rate of 0.001.
-For the testing stage, we freeze all the parameters of our backbone to extract task-specific discriminative video features and semantic features for measuring the distances.
+We freeze the original pre-trained model and only fine-tune the introduced Task adapters and order adapters using the SGD optimizer with a learning rate of 0.001.
 All reported results in our paper are the average accuracy over 10,000 few-shot action recognition tasks.
-Table 2: This part reported the results focused on scene-related datasets (including HMDB51, UCF101, and Kinetics).
-Comparison with current SOTA few-shot action recognition methods on 5-way 1-shot, 5-way 3-shot and 5-way 5-shot benchmarks.
+Table 2: Comparison with current SOTA few-shot action recognition methods on scene-related datasets (HMDB51, UCF101 and Kinetics) with 5-way 1-shot, 5-way 3-shot and 5-way 5-shot settings.
 The best results are highlighted in bold.
@@
 HyRSM++ [58] PR’24 ResNet50 IN-21K Full Fine-tuning 61.5 72.7 76.4 85.8 93.5 95.9 74.0 83.9 86.4
-[…]
+MASTAF [59] WACV’23 ViViT JFT Full Fine-tuning 69.5 - - 91.6 - - - - -
 SA-CT [49] MM’23 ViT-B IN-21K Full Fine-tuning - - 81.6 - - 98.0 - - 91.2
@@
 Task-Adapter MM’24 ViT-B CLIP PEFT 83.6 88.0 88.8 98.0 98.7 99.4 95.0 96.3 96.8
-Task-Adapter++ - ViT-B CLIP PEFT 86.5 90.6 92.0 98.9 99.5 99.7 96.0 97.3 97.5
+Task-Adapter++ - ViT-B CLIP PEFT 86.5 90.6 92.0 98.9 99.5 99.7 96.2 97.3 97.5
 4.2 Comparison with state-of-the-art methods
-Table 1 and Table 2 presents a comprehensive comparison of our method with recent few-shot action recognition methods.
+Table 1 and Table 2 presents a comprehensive comparison of our method with recent few-shot action recognition methods on both temporal-related datasets and scene-related datasets.
 As our Task-Adapter++ is designed for ViT-based architectures, we use CLIP [1] pre-trained ViT to make a fair comparison with recent few-shot action recognition works that use large-scale pretrained models.
 Of particular note, by comparing the results of works that use CLIP pre-trained models and those that use ImageNet (IN-21K) pre-trained models, we can conclude that the performance of few-shot action recognition indeed benefits from large-scale pretraining, especially for scene-related datasets.
-As shown in Table 1, our Task-Adapter++ can achieve best results on SSv2-Small and SSv2-Full, indicating that our method can model the temporal information well.
-For Average 1-shot, 3-shot and 5-shot, our method outperforms the exsisting SOTA methods by 3.1% on SSv2-Small and 1.4% on SSv2-Full, respectively.
-Compared with full fine-tuning method CLIP-FSAR [41], we can see our method has an overall improvement (from 3.4% to 11.4%).
-Compared with PEFT method TSAM [17], our method also has an overrall improvement (from 0.9% to 7.0%) on two datasets.
-We generally believe that SSv2 requires more temporal modeling than other three datasets, making it much more challenging, and we have achieved a significant improvement of at least 3% on SSv2-Small compared to the SOTA, indicating that our task can better adapt to the classification of complex actions.
-In addition, we display the results on HMDB51, UCF101 and Kinetics in Table 2.
-It is obvious that we achieve SOTA results than other methods.
-Compared with full fine-tuning method CLIP-FSAR [41], we achieve an average improvement of 6.9% on HMDB51, 1.3% on UCF101 and 1.7% on Kinetics.
-Compared with PEFT method TSAM [17], our method also has an average improvement of 2.6% on HMDB51, 0.5% on UCF101 and 0.1% on Kinetics.
+As shown in Table 1, our Task-Adapter++ can achieve the best results on SSv2-Small and SSv2-Full, leading previous SOTA method with a large margin, which indicates the superiority of our method on temporal challenging few-shot learning tasks.
+The results of Task-Adapter compared with MA-FSAR [19] on the SSv2 datasets demonstrate the effectiveness of our proposed method for task-specific adaptation, showing a significant improvement of 5.7% for the 5-shot task on SSv2-Small and 8.0% for the 1-shot task on SSv2-Full, respectively.
+Our Task-Adapter have achieved an average improvement of 3.4% on SSv2-Small and 4.9% on SSv2-Full compared to the SOTA method MA-FSAR [19].
+Our Task-Adapter++ have achieved further improvement on the basis of Task-Adapter, specifically improving 3.1% on SSv2-Small and 1.4% on SSv2-Full, benefiting from the semantic order adaptation and fine-grained cross-modal matching ability.
+Compared with full fine-tuning method CLIP-FSAR [60], we can see our Task-Adapter++ has a significant overall improvement (from 3.4% to 11.4%) with fewer learnable parameters (7.5 M vs.
+89.3 M).
+Compared with PEFT method TSAM [21], our Task-Adapter++ also has a comprehensive improvement (from 0.9% to 7.0%) on SSv2.
+Our Task-Adapter++ has further expanded the average improvement, 4.3% on SSv2-Small and 3.9% on SSv2-Full compared to the SOTA method TSAM [21], indicating that our method can better adapt to the temporal complex FSAR tasks with dual adaptation.
+In addition, we report the results on HMDB51, UCF101 and Kinetics in Table 2.
+Especially for 5-shot tasks, our Task-Adapter outperforms the existing SOTA methods MA-FSAR [19] by 0.9% on HMDB51, 0.2% on UCF101 and 0.8% on Kinetics, respectively.
+Above our Task-Adapter, Task-Adapter++ achieve higher score by 2.9% on HMDB51, 0.7% on UCF101 and 1.0% on Kinetics, respectively.
+Compared with full fine-tuning method CLIP-FSAR [60], our Task-Adapter++ achieve an average improvement of 6.9% on HMDB51, 1.3% on UCF101 and 1.8% on Kinetics.
+Compared with PEFT method TSAM [21], our method also has an average improvement of 2.6% on HMDB51, 0.5% on UCF101 and 0.2% on Kinetics.
 These scene-related datasets require strong background understand ability which means our model can handle complex scene knowledge well.
-Table 3: We choose frozen backbone as the baseline and compare the performance of different settings for 5-way 1-shot task on SSv2-Small.
+Table 3: Ablation study of Task-adapter. We choose frozen backbone as the baseline under 5-way 1-shot settings on SSv2-Small.
 Frozen Backbone AIM Task-specific Adaptation Partial Adapting Acc
@@
 $\times$ $\times$ 53.3
-[…]
+$\times$ 54.3
 $\times$ 55.2
@@
 60.2
-Table 4: We choose our Task-Adapter as the baseline and compare the performance of different settings for 5-way 1-shot task on SSv2-Small.
-Task-Adapter Three Sub-actions Order Adaptation Cross Attention Alignment Acc
+Table 4: Ablation study of Task-adapter++. Task-Adapter is chosen as the baseline under 5-way 1-shot settings on SSv2-Small.
+Task-Adapter Three Sub-actions Order Adaptation Cross-modal Alignment Acc
 $\times$ $\times$ $\times$ 60.2
@@
 $\times$ $\times$ 58.5
-[…]
+$\times$ 62.5
 $\times$ 62.8
@@
 Effectiveness of each component.
-To demonstrate the effectiveness of each component, we perform detailed ablation experiments in Table 3 and Table 4.
+To demonstrate the effectiveness of each component, we perform detailed ablation experiments.
+The ablation study of our Task-Adapter is show in Table 3 and Task-Adapter++ is shown in Table 4.
 In Table 3, the first row shows the results of a frozen pre-trained model without any learnable adapters, while the second row refers to the performance of full fine-tuning.
 Comparing the results of the third and fourth rows, we can see the effectiveness of PEFT over traditional full fine-tuning in few-shot action recognition.
-Note that the "Partial Adapting" in the fourth column of Table 3 refers to a common strategy which only adds adapters in the last several ViT layers.
+Note that the “Partial Adapting” in the fourth column of Table 3 refers to a common strategy which only adds adapters in the last several ViT layers.
 We can observe from the table that the improvement of introducing task-specific adaptation into AIM is distinct (1.9%) when inserting adapters into all the layers of the pre-trained model.
@@
 Even so, the improvement of introducing task-specific adaptation into AIM is notably enhanced under the partial adapting strategy, increasing from 1.9% to 5.9%.
-In Table 4, the first row presents the baseline result, while the second row shows the performance when applying the three sub-actions, which yields a 1.8% improvement.
-The third row reports the effect of applying cross attention alignment, leading to a slight decline of 1.7%.
-Notably, performing an alignment operation without enhancing semantic features is redundant, as the original Task-Adapter already accomplishes semantic alignment during feature extraction.
-This observation in the fourth row supports our assumption, where the model’s performance increases further from 1.8% to 2.3%.
-Incorporating order adaptation improves the performance of the Three Sub-actions.
+In Table 4, the first row presents the result of Task-Adapter, while the second row shows the performance when applying the three sub-actions for textual augmentation, which yields a 1.8% improvement.
+The third row reports the effect of applying cross-modal alignment with fine-grained temporal modeling, leading to a decline of 1.7%.
+Notably, performing the fine-grained alignment operation without enhancing semantic features is redundant, as the original Task-Adapter already accomplishes semantic alignment during feature extraction.
+The result in the fourth row supports our assumption, where the model’s performance increases further from 1.8% to 2.3%.
+Incorporating order adaptation improves the performance of the “Three Sub-actions”.
 It models the sequential relationships between sub-actions in more detail.
-This change increases performance from 1.8% to 2.5% as shown in the fifth row.
-The final row summarizes the complete configuration.
-Figure 8: Effect of inserting Task-Adapters into the last $L$ ViT layers (e.g., $L$ = 1, 2, 3, 6, 12) on scene-related datasets. The best is highlighted with a red star.
-Figure 9: Effect of inserting semantic order adapters into the last $M$ ViT layers (e.g., $M$ = 1, 2, 3, …, 12) on 3 different datasets. The best is highlighted with a black star.
+This change increases the performance from 1.8% to 2.5% as shown in the fifth row.
+The last row represents the performance of the complete model, revealing the complementarity and effectiveness of each component.
+Figure 8: Effect of inserting Task-Adapter into the last $L$ layers (a) and semantic order adapter (b) into the last $M$ layers (e.g., $L,M$ = 1, 2, 3, …, 12) on 3 different datasets.
+The best is highlighted with a black star.
+The effect of inserting position of adapters.
+In this part, we explore the result of different inserting positions.
+We can see from Fig.
+8 (a) that inserting our Task-Adapter into the top 2 layers is the best setting for Kinetics and HMDB51.
+Since these scene-related datasets focus on background understanding, a more lightweight setting with inserting adapters into only the top 2 layers is more appropriate.
+Inserting the Task-Adapter into the top 6 layers yields the highest performance on SSv2.
+Since SSv2 requires more intensive temporal modeling than scene-related datasets, a deeper adapter is necessary.
+In Fig.
+8 (b), we demonstrate the impact of the insertion position of the semantic order adapter.
+Our analysis indicates that for datasets such as SSv2—which already encode rich action information in their original labels (e.g., “Pouring something into something”)—employing a deep architecture for text sequence modeling is generally unnecessary.
+Conversely, for datasets with more straightforward action content, such as Kinetics and HMDB51, a more sophisticated modeling approach is warranted.
+In conclusion, for scene-related datasets like UCF101, Kinetics and HMDB51, we choose 2-layer Task-Adapter and 8-layer semantic order adapter.
+For SSv2, we choose 6-layer Task-Adapter and 2-layer semantic order adapter.
+A detailed analysis of the tunable parameters will be presented in the following subsection.
 Table 5: The result that using different modalities of information.
-methods visual only semantic only both
+methods video matching score video-text matching score both
 Task-Adapter 53.4 50.4 60.2
 Task-Adapter++ 53.1 54.3 63.6
-Figure 10: We demonstrate accuracy for each time step computed by a single sub-action description to prove its rationality.
+The result of different modalities of information.
+As shown in Table 5, the integration of multimodal information into the FSAR task significantly enhances the performance of both Task-Adapter and Task-Adapter++.
+However, we can observe that Task-Adapter achieves an accuracy of only 50.4% when relying solely on the video-text matching score, suggesting that semantic information is not effectively leveraged.
+While, Task-Adapter++ can significantly improve the classification accuracy of the semantic branch by 3.9%.
+This improvement can be attributed to the proposed order adaptation of semantic information and fine-grained cross-modal matching.
+Our method effectively addresses the lack of informative cues in video-text matching, balancing the two branches of the model and enhancing their complementary strengths.
+Table 6: The results of different utilization methods for enhanced sub-action description.
+methods w/o sub-actions average features concat sub-actions ours
+accuracy 60.2 60.7 61.4 63.6
+Different methods for using the sub-actions.
+In general, LLMs provides three descriptions of sub-actions.
+Except for our claimed method in Section 3, it is easy to consider them as a whole.
+This actually includes two ideas: averaging the three semantic features or concatenating the descriptions together as one prompt.
+Therefore, Table 6 shows the accuracy with different strategies, i.e., without using sub-actions description, averaging features and concatenating descriptions together.
+The results not only demonstrate the effectiveness of decomposing actions into sub-actions, but also reveal the information loss caused by averaging and concatenation, thereby validating the rationale behind stage-wise matching
 Figure 9: The results of $N$-way 1-shot between different methods on SSv2-Small and Kinetics.
-The effect of inserting position of adapters.
-We can see from the Fig.
-8 that inserting our Task-Adapter into the top 2 layers is the best setting for all these three datasets.
-Since these scene-related datasets focus on background understanding and do not need much temporal modeling, a more lightweight setting with inserting adapters into only 2 top layers is more appropriate.
-In the Fig.
-9, we demonstrate the impact of the insertion position of a semantic order adapter.
-Our analysis indicates that for datasets such as SSv2, which already encode rich action information within their original labels, employing a deep architecture for text sequence modeling is generally unnecessary.
-We posit that the labels in SSv2, characterized by an abundance of verbs and positional terms, combined with the enhanced contextual information from LLM, yield comprehensive representations.
-Conversely, for datasets with more straightforward action content, such as Kinetics and HMDB51, a more sophisticated modeling approach is warranted.
-The result that using different modalities of information on SSv2-Small 5-way 1-shot.
-As shown in Table 5, we observe that Task-Adapter++ can significantly improve the classification accuracy of the semantic branch by 3.9%.
-This improvement is undoubtedly due to the proposed augmentations of semantic information.
-Our method effectively fills the gap of insufficient information in semantic visual matching, making the two legs of the model equally long.
-The accuracy for each frames.
-In general, LLM provides three descriptions of sub-actions.
-Except for our claimed method in Section 3, it is easy to consider them as a whole.
-This actually includes two ideas: averaging three semantic features or concatenating the descriptions together as one prompt.
-Therefore, Fig.
-10 shows the accuracy with different baselines, i.e., without using sub-actions description, averaging features and concatenating three descriptions together.
-Following Section 3, we separately compute accuracy for each temporal moment using the 3 sub-action descriptions.
-The results not only indicate that describe sub-actions in this manner is reasonable and helpful, but explain the information loss when just averaging or concatenating them.
 $N$-way 1-shot results.
 Fig.
-11 shows that our approach, compared to the original Task-Adapter, not only improves absolute accuracy but also better adapts to $N$-way 1-shot learning ($N$geq 5), i.e., experiences less relative decline.
+9 shows that our approach, compared to the original Task-Adapter and the baseline, not only improves absolute accuracy but also better adapts to $N$-way 1-shot learning ($N$geq 5), i.e., experiences less relative decline.
 As $N$ increases, classification difficulty also increases, and model performance decline is inevitable.
-However, because our approach fully considers both semantic and visual features, the model will be less affected in SSv2-Small.
+However, because our approach fully considers both semantic and visual features, its performance is less affected in SSv2-Small.
 And in Kinetics, the results are nearly saturated.
 It can be seen that the results descend slightly when $N$ increases, and the gap keeps very small.
-Figure 12: The choice of different alignment modules on SSv2-Small 5-way 1-shot.
+Table 7: The results of different alignment modules on 5-way 1-shot task.
+Method setting SSv2-Small Kinetics
+1-shot 63.6 96.2
+1-layer Cross Attnetion 3-shot 69.1 97.3
+5-shot 73.2 97.5
+1-shot 60.4 94.3
+2-layer Cross Attnetion 3-shot 66.7 96.1
+5-shot 71.7 97.1
+1-shot 60.6 93.3
+3-layer Cross Attnetion 3-shot 67.0 95.7
+5-shot 71.4 96.8
+1-shot 55.4 92.3
+Conv1d 3-shot 62.9 94.9
+5-shot 68.5 96.0
 The exploration of different alignment modules.
-In Fig.
-12, we explore different modules for the alignment task.
+In Table 7, we explore different temporal modeling modules for the alignment task.
 We utilize multi-layer cross attention and 1D CNN (layer=1, stride=1, kernel=3, channel=512) to align visual features.
-The results indicate that stacking multi-layer cross attention is ineffective and does not improve results, while 1D CNN shows no advantages over cross attention.
+The results indicate that stacking multi-layer cross attention is ineffective and does not improve results, while 1D CNN shows significant disadvantage over cross attention.
 Figure 10: Comparison of the performance achieved by combining different fine-tuning strategies with existing widely used metric measurement methods on SSv2-Small 5-way 1-shot.
@@
 Taking full fine-tuning strategy as the baseline, we observe a consistent improvement across all metric measurement methods, as shown in Fig.
-13.
+10.
 It should be noted that ProtoNet outperforms all other methods when uniformly using Task-Adapter++.
-We attribute this to the fact that with our Task-Adapter++, the model significantly improves the utilization rate of multimodal information and enhances fault tolerance on the basis of Task-Adapter.
-Table 6: Effect of inserting position of task adapters and order adapters on SSv2-Small.
+We attribute this to the fact that with our Task-Adapter++, the model significantly improves the utilization of multimodal information and enhances fault tolerance on the basis of Task-Adapter.
+Table 8: The tunable parameter analysis of task adapters and order adapters on SSv2-Small.
 Task-Adapter Order Adapter
@@
 Bottom 6 7.2 M 50.7 Bottom 6 790 K 62.1
-Bottom 2 2.4 M xxx Bottom 2 263 K 62.6
+Bottom 2 2.4 M 47.4 Bottom 2 263 K 62.6
 Top 6 7.2 M 60.2 Top 6 790 K 61.2
 Top 2 2.4 M 54.3 Top 2 263 K 63.6
-Dual Partial Adapting.
-Firstly, we study the effect of different inserting positions of our Task-Adapter in Table 6 left.
-The default setting is inserting adapters into all of the 12 ViT layers.
-Alternatively, we also evaluate the settings which respectively insert adapters to the bottom 2/6 layers and top 2/6 layers of the model (assuming forward propagation is from bottom to top).
-Inserting adapters only into the top 6 layers has been shown to achieve comparable performance with inserting to all layers in AIM [54].
-However, inserting our Task-Adapter only into the top 6 layers outperforms inserting it into all layers by a large margin (5%) for few-shot action recognition.
-We attribute this to the hypothesize that bottom layers focus on generic features that do not require extensive adaptation, while the top layers focus on task-specific and discriminative features that benefit a lot from our task-specific adaptation module.
-Additionally, We study the effect of different inserting positions of our semantic order adapter in Table 6 right.
-The default setting is inserting adapters into all of the 12 ViT layers.
-We evaluate settings inserting adapters to bottom 2/6 and top 2/6 layers.
-Notice that inserting adapters only top 2 layers achieves best performance with only 263 K parameters.
-Considering little difference in results at bottom 2/6 layers, we believe the semantic features have not been well extracted yet, which is exactly the opposite at the top 2/6 layers.
+The tunable parameter analysis.
+Firstly, we study the accuracy and tunable parameter of different inserting positions of our Task-Adapter in Table 8 left.
+On SSv2-Small, inserting our Task-Adapter only into the top 6 layers outperforms inserting it into all layers by a large margin (5%) with only 7.2 M parameters.
+We attribute this to the hypothesize that bottom layers focus on generic features that do not require extensive adaptation, while the top layers focus on task-specific and discriminative features that benefit a lot from task-specific adaptation module.
+Additionally, We study the accuracy and tunable parameter of different inserting positions of our semantic order adapter in Table 8 right.
+Note that inserting adapters only into the top 2 layers achieves the best performance with only 263 K parameters.
+Given the small performance gap at the bottom 2/6 layers, we believe that semantic features are not yet well extracted, which is in contrast to the top 2/6 layers where stronger semantic representations are observed.
+In conclusion, for scene-related datasets, Task-Adapter++ utilizes only 2.4 M parameters in the visual branch and 1.1 M parameters in the semantic branch, resulting in a total of 3.5 M parameters.
+In comparison, for temporal-related datasets, the model employs 7.2 M parameters in the visual branch and 263 K parameters in the semantic branch, amounting to 7.5 M parameters in total.
+These two strategies both significantly reduce the number of parameters compared to full fine-tuning, which requires 149.6 M parameters.
 Table 9: Impact of position of Task-MSA and O-MSA inner one ViT layer on SSv2-Small 5-way 1-shot.
@@
 Position of the Task-MSA and O-MSA inner ViT layer.
-In the visual branch, as illustrated in Section 3, the original AIM comprises two multi-head self-attention (MSA) modules, individually for spatial and temporal modeling, and a task adapter.
-We study the optimal placement of our proposed task-specific MSA (Task-MSA) relative to the existing two MSAs for improved task modeling.
-As shown in Table  7 left, inserting the Task-MSA after the temporal (T-) and spatial (S-) MSAs achieves the best performance.
+In the visual branch, as illustrated in Section 3, the original AIM comprises two multi-head self-attention (MSA) modules, individually for spatial and temporal modeling.
+We study the optimal placement of our proposed task-specific MSA (Task-MSA) relative to the existing two MSAs for task-specific modeling.
+As shown in Table 9 left, inserting the Task-MSA after the temporal (T-) and spatial (S-) MSAs achieves the best performance.
 It is noteworthy that the Task-MSA consistently outperforms the baseline AIM [5], regardless of its insertion position.
@@
 Furthermore, placing the Task-MSA after the T- and S-MSAs enables maximum extraction of task-specific discriminative features.
-In the semantic branch, we retain the original Task-Adapter as our baseline without changes to the encoder.
-We find that relocating the order MSA (O-MSA) back of the original MSA yields higher scores than placing it frontally, as shown in Table 7 right.
-This is because we model the order relationship of sub-action sequences at the feature level, so positioning the O-MSA later enables better feature-level modeling.
-Nonetheless, both placements perform better than the baseline Task-Adapter, proving the effectiveness of the Order-MSA.
+In the semantic branch, we retain the original Task-Adapter as our image encoder.
+We find that placing the order MSA (O-MSA) after the original MSA yields higher scores than placing it frontally, as shown in Table 9 right.
+This is probably because modeling the order relationships of sub-action sequences benefits more from refined representations, which are typically captured in later stages of the network.
+Nevertheless, both configurations outperform the baseline Task-Adapter, demonstrating the effectiveness of the proposed O-MSA
+Figure 11: Visualization of the attention maps and cross-modal matching scores for the “Pouring something into something” action of different methods.
+We first present the attention maps generated by AIM and Task-Adapter++ to demonstrate the superiority of our Task-Adapter.
+Subsequently, we illustrate the cross-modal matching scores at various stages, thereby substantiating the effectiveness of the dynamic sub-action description.
+4.4 Visualization results
+To further demonstrate the effectiveness of our proposed Task-Adapter++, we visualize the attention maps and cross-modal matching scores in Fig.
+11.
+The input frames are sampled from the “Pouring something into something” action in SSv2-Small.
+Notably, the attention map generated by AIM is less effective at capturing informative regions, such as the carton of milk, compared to that produced by our Task-Adapter++.
+Furthermore, our method employs a dynamic sub-action description that segments the complete action into three distinct stages—beginning, process, and end.
+In comparison with the static prompt, our dynamic prompting strategy yields significant improvements in fine-grained cross-modal matching performance.
 5 Conclusion
-In this paper, we propose a novel method named Task-Adapter++ to better adapt the pre-trained image models for few-shot action recognition.
-Instead of using the traditional full fine-tuning strategy, we propose a parameter-efficient dual adaptation method that only fine-tune the newly inserted adapters which alleviates the issues of catastrophic forgetting and overfitting.
-Specifically, we design task-adapter module to enhance the most discriminative features within the given few-shot learning tasks in visual branch.
-We otherwise utilize semantic order adapter to model the sequence of sub-actions in semantic branch.
-We further apply cross attention for cross modal alignment operations, which remarkably improves the performance on five few-shot action recognition benchmarks and achieves the new state-of-the-art results, particularly on the challenging SSv2 dataset.
+In this paper, we propose Task-Adapter++ , a novel framework for adapting pre-trained image models to few-shot action recognition.
+Instead of traditional full fine-tuning, we introduce a parameter-efficient dual adaptation method that exclusively fine-tunes newly inserted adapters, thereby mitigating catastrophic forgetting and overfitting issues.
+Specifically, we design a task adapter module to enhance discriminative features in the visual domain for given few-shot learning tasks.
+Concurrently, we generate sub-action descriptions to capture dynamic label information and employ a semantic order adapter to model the sequential structure of sub-actions in the semantic domain.
+To better align visual and semantic features for cross-modal matching, we further propose a novel fine-grained cross-modal alignment strategy.
+Experimental results fully demonstrate the effectiveness and superiority of the the proposed Task-Adapter++, which yields significant performance improvements across five few-shot action recognition benchmarks, particularly on the challenging SSv2 dataset.
 6 Acknowledgements
-xxx
+This work is supported by National Natural Science Foundation of China (No.
+62376217, 62301434), Young Elite Scientists Sponsorship Program by CAST (No.
+2023QNRC001), and the Joint Research Project for Meteorological Capacity Improvement (Grants 24NLTSZ003).